### v2

Abstract
People, robots, and companies mostly divide time and effort between projects, and shared effort games model people investing resources in public endeavors and sharing the generated values. In linear $\theta$ sharing (effort) games, a project’s value is linear in the total contribution, thus modelling predictable, uniform, and scalable activities. The threshold $\theta$ for effort defines which contributors win and receive their share, equal share modelling standard salaries, equity-minded projects, etc. Thresholds between 0 and 1 model games such as paper co-authorship and shared assignments, where a minimum positive contribution is required for sharing in the value. We constructively characterise the conditions for the existence of a pure equilibrium for $\theta\in\{0,1\}$, and for two-player games with a general threshold, and find the prices of anarchy and stability. We also provide existence and efficiency results for more than two players, and use generalised fictitious play simulations to show when a pure equilibrium exists and what its efficiency is. We propose a novel method for studying solution concepts by defining a new concept and proving its equivalence to a previously known on a large subclass of games. This means that the original concept narrows down to a more demanding concepts on certain games, providing new insights and opening a path to study both concepts conveniently. We also prove mixed equilibria always exist and bound their efficiency.
keywords:
Shared effort games , Contribution threshold , Equilibrium , Price of anarchy , Price of stability , fictitious play , narrowing down/broadening (a solution concept) , cyclically strong equilibrium
journal: Discrete Applied Mathematics
[…]
[label5]organization=Delft University of Technology, addressline=Van Mourik Broekmanweg 6, city=Delft, postcode=2628XE, country=The Netherlands
1 Introduction
Many real-world situations include a group of players investing resources across multiple projects. Examples of such situations include financial investments, contributions to online communities [21], Wikipedia [16], political campaigns [49], paper co-authorship [24], social exchange networks [25]. Naturally, also hobbies and attention spread fall under this category. In the formal analyses of these type of games, it is often assumed that the obtained revenue from such projects is linear in the total contribution and is shared equally [44, 24, 8]. Basically, performing standard predictable activities, which are scalable and homogeneous, often yields linear value (and we also tend to linear thinking [15]). Equal division takes place with standard salaries, fundraising for community projects, equal distribution of recognition to all the contributors, equity-minded projects (the same reward regardless the contribution), non-rivalrous values, like public parks, which can be enjoyed equally by all those who have access to them, etc. However, in most of the above examples, revenue is only shared among those who contribute at least a certain amount of effort. In this work we model this as a threshold on a player’s contribution. In particular, we analyse the situation when the considered threshold is relative to the investments of the other players. The abstract model of this paper is the first theoretical study of such shared effort games with a threshold, assuming linear project values. This provides the foundations for modelling contributions in various kinds of projects, basically parallel games, such as considering splitting time between projects, each of which is a reciprocal interaction [41].
An example with a relative threshold of $1$ is an all pay-auction. Another example is the Colonel Blotto game (see e.g. [44]), where there are several battlefields over which each of the $2$ players distributes her soldiers, and the winner at each battlefield is the one who has put more people there. A player’s payoff is the number of projects (battlefields) where she has won, the value of each project being fixed. Thus, when our model is reduced to $2$ players with a fixed value of each project, we obtain the Colonel Blotto competition, whereas we typically assume a project has a value that increases as a function of the total contribution, modelling constructive collaboration, rather than countenance. These examples are “highly thresholded” because only the player whose effort per project is maximum collects the complete revenue.
We now present examples of roughly linear growing project values, each being equally shared between all those who contribute at least a certain threshold.
  1. 1.
    Assigning points for an exercise, where a percentage of the perfect work is required to obtain the (equal) homework’s credits [20] is an absolute threshold example from a course at the University of Maryland. Linearity is a reasonable model for predictable tasks. Although our threshold is relative, while in the Maryland example it is absolute, still, professors often relate the individual grades to the average level of the students.
  2. 2.
    Consider kids selling lottery tickets for their sports club where a significant part of the income is used to improve the club’s facilities; kids that sell at least a certain number (threshold) of tickets, relatively to the others, are equally rewarded by the club with the same symbolic present (e.g., club memorabilia, like a hat, a scarf, or a jersey).
  3. 3.
    Another example is a start-up where developers are building a new tool to sell and divide the revenue equally, or any firm paying equal salaries to everyone who works enough to keep the job. Here, the resulting value is assumed to grow approximately linearly in the contributed time, when the idea being implemented is clear and no surprises are expected.
The following example with a smaller threshold is used later to further illustrate the model.
Example 1.
Consider two collaborating scientists in a narrow research field. They can work on their papers alone or together. When they collaborate on a paper (a project), an author has to contribute at least $0.2$ of the work of the other one, in order to be considered a co-author. The value of a paper, being the recognition, is equally divided among the authors. Author $1$ has the time budget of $5$ hours to work, and author $2$ has $20$ hours. The value of the first paper is $4$ times the total time the authors put in, while the less “hot” second paper rewards the contributors with a value of only twice the contributed time. This is illustrated in Figure 1. In the figure, the first paper receives the total contribution of $4+10=14$, creating the value of $4\cdot 14=56$. Both contributors are authors, since $4\geq 0.2\cdot 10$, and the value is equally divided between them. The second paper receives $1+10=11$, and yields the value of $2\cdot 11=22$. Here, only the contributor of $10$ is an author, since $1<0.2\cdot 10$, and he, thus, receives the whole value of $22$. This is not a Nash equilibrium, since the second contributor would benefit from moving the $1$ hour contribution to the first paper, increasing her share from the first paper by 2. On the other hand, if both authors invest all their time in paper $1$, the situation is stable. Indeed, moving a part to paper $2$ would benefit nobody, since the paper is twice less valuable than paper $1$, so sharing the value of paper $1$ is as good as contributing alone to paper $2$. The social welfare in this equilibrium is maximum possible, since everyone contributes to the most valuable project. In general, we would like to find stable contributions, and whether they will be efficient for both authors, relatively to the maximum possible divisions of the authors’ time budgets.
[…]
Figure 1: The co-authors invest what is shown in the arrows that go up, every project’s value is defined as the $P$ function of the total contribution, and it is equally shared among the contributors who contribute above the relative threshold of $0.2$. The obtained shares are denoted by the arrows that go down.
It is important to understand and predict the stable contribution strategies, and suggest the efficient ones in situations where players invest resources in several projects and share the generated values. The classical stability model is the Nash equilibria, and the efficiency model of stable situations is the ratio of the least total utility of the players in an equilibrium to the optimum, called the price of anarchy ($\operatorname{PoA}$[27], and the ratio of the largest total utility in equilibrium to the optimal total utility, called the price of stability ($\operatorname{PoS}$[47, 4]. If the price of anarchy is close to $1$, then all equilibria are good, and we may follow any equilibrium profile. If the price of anarchy is low, while the price of stability is high, then we have to regulate the play by suggesting the efficient equilibria, while if even the price of stability is low, the only way to make the play socially efficient is changing the game through, for example, subsidising.
This price of anarchy for shared effort games was bounded in [8], but assuming that a player obtains at least $1/k$th of her marginal contribution, which fails to hold in a positively thresholded model. Threshold was introduced in [43], and pure equilibria for two players with close budgets and projects with linear utility functions were characterised. There is no analysis of the existence of Nash Equilibria (NE) and their efficiency in general linear shared effort games and nothing about mixed Nash equilibria is known. This paper fills this gap with the following contributions. [^4] Relatively to [43], we extend the theory also for budgets not within a threshold factor from each other, characterise the non-suppressed equilibria for threshold $1$, provide the novel concept of narrowing down Nash to the newly defined cyclically strong equilibrium, prove the existence of an NE in the mixed case, answering this natural question, and substantially extend the simulations.
Our Contributions
For $2$ players,
a complete constructive characterization of the existence of pure NE in Theorem 3, and exact values of the price of anarchy and stability.
For more players,
analysing the cases of 00 threshold and $1$ threshold (Theorem 2), sufficiency results on existence of pure NE and efficiency bounds, and exploring this domain through simulation of fictitious play to find pure NE and their efficiency in 2-project multi-player games. The simulations need to be engineered based on the case in question, so the importance of the simulations is mainly in the proof of concept of implementing the infinite strategy fictitious play for 2 projects. This involves designing an $O(n\log n)$ best response algorithm for $2$-project multi-player games and proving that for more projects, best-responding is NP-hard, even for the same project functions and $2$ players.
For any number of players,
we introduce the concept of narrowing down (or broadening) a solution concept. In our case, we strengthen Nash equilibrium by adding a requirement of no profitable cyclical deviations of whole budgets, alluding to the famous top trading cycles mechanisms [1]. We then demonstrate that when no budget can make another one fall below the threshold, both solution concepts coincide.
For any number of players,
a proof of existence of mixed NE (it is an infinite game, so Nash’s Theorem [38] does not apply) and extending efficiency bounds to mixed equilibria.
After defining shared effort games in the next section, we first concentrate on pure strategies. We theoretically treat the existence and efficiency of NE for thresholds 00 and $1$ and then for games with two players in Section 3. Afterwards, we study games with any number of players. In Section 4, we define a new solution concept and narrow down Nash equilibria to that concept, whetting the appetite for defining further solution concepts and studying their relationships by analogy. Finally, we treat the mixed extension in Section 5, proving the existence of a mixed NE and showing some efficiency bounds of the pure case generalise to the mixed case. We conclude and discuss the future work in Section 7. When a proof is mostly technical, we provide its sketch and intuition in the body of the paper and defer the complete proof to A. We conduct further investigation of existence and efficiency of pure NE for more than $2$ players by employing fictitious play simulations in B. Concretely, we define an Infinite-Strategy Fictitious Play and simulate it by computing a best response, till and if it finds an equilibrium within some time. We check whether we have found an NE and if that is the case, what its efficiency is.
We introduce and study the threshold on the basic model, motivating more refinements of the model. This is a major impact, alongside the mothodological implications of the notion of narrowing down of equilibria from Section 4, which motivates defining other appropriate solution concepts and seeing when a narrowing down or a broadening occurs.
2 Model
To model investing effort in shared projects, we define shared effort games, extending the model by [8] with thresholds. The games consist of players who contribute to project, and share the value the projects generates. Henceforth, we use the general term “value” instead of the monetary “revenue”. Formally, players $N=\left\{1,\ldots,n\right\}$ contribute to projects $\Omega$. The contribution of player $i$ to project $\omega\in\Omega$ is denoted by $x^{i}_{\omega}\in{\mathbb{R}_{+}}$. Each player $i$ has a budget $B_{i}>0$, and the strategy space of player $i$ (i.e., the set of her possible actions) is [^5] We denote “defined as” by =$\stackrel{{\scriptstyle\rm\Delta}}{{=}}$. $S^{i}\stackrel{{\scriptstyle\rm\Delta}}{{=}}\left\{x^{i}=(x^{i}_{\omega})_{\omega\in\Omega}\in\mathbb{R}_{+}^{\left|\Omega\right|}\mid\sum_{\omega\in\Omega}{x^{i}_{\omega}}\leq B_{i}\right\}$. Assume w.l.o.g. (without loss of generality) that $B_{n}\geq\ldots\geq B_{2}\geq B_{1}>0$. Being an $\left|\Omega\right|-1$-dimensional simplex, this set is compact and convex. Denote the vector of all the contributions by $x=(x^{i}_{\omega})^{i\in N}_{\omega\in\Omega}$ and the strategies of all the players except $i$ by $x^{-i}$.
To define the utilities, each project $\omega\in\Omega$ is associated with its (really or perceivably) linear (allows concentrating on cumulative scalable cooperation across multiple “projects”, like manufacturing, technical writing, Wikipedia or environmental activism) project function, which determines its value, based on the total contribution $x_{\omega}\stackrel{{\scriptstyle\rm\Delta}}{{=}}\sum_{i\in N}{(x_{\omega}^{i})}$ that the project receives; formally, $P_{\omega}(x_{\omega})\colon{\mathbb{R}_{+}}\to{\mathbb{R}_{+}}$, and $P_{j}(y)=\alpha_{j}y$. W.l.o.g., $\alpha_{m}\geq\alpha_{m-1}\geq\ldots\geq\alpha_{1}>0$. We denote the number of projects with the largest coefficient of project functions by $k\in\mathbb{N}$, i.e. $\alpha_{m}=\alpha_{m-1}=\ldots=\alpha_{m-k+1}>\alpha_{m-k}\geq\alpha_{m-k-1}\geq\ldots\geq\alpha_{1}$. We call those $k$ projects steep. The project’s value is distributed among the players according to the following rule. From each project $\omega\in\Omega$, each player $i$ that contributes enough gets an equal share, denoted $\phi_{\omega}^{i}(x_{\omega})\colon{\mathbb{R}_{+}}\to{\mathbb{R}_{+}}$. Such sharing is relevant to many applications where a minimum contribution is required to share the value, such as paper co-authorship and homeworks, and we study predominantly such games. For any $\theta\in[0,1]$, define the players who get a share as those who bid at least a $\theta$ fraction of the maximum bid size to $\omega$,
$N^{\theta}_{\omega}\stackrel{{\scriptstyle\rm\Delta}}{{=}}\left\{i\in N_{\omega}\mid x^{i}_{\omega}\geq\theta\cdot\max_{j\in N_{\omega}}{x^{j}_{\omega}}\right\}.$
The $\theta$-equal sharing mechanism equally divides the project’s value between all the users who contribute at least $\theta$ of the maximum bid to the project, as we now formally define.
Definition 1.
The $\theta$-equal sharing mechanism, denoted by $M_{\text{eq}}^{\theta}$, is
$\phi^{i}_{\omega}(x_{\omega})\stackrel{{\scriptstyle\rm\Delta}}{{=}}\begin{cases}\frac{P_{\omega}(x_{\omega})}{\left|N^{\theta}_{\omega}\right|}&\mbox{if}i\in N^{\theta}_{\omega},\\ 0&\mbox{otherwise}.\\ \end{cases}$
The utility of a player $i\in N$ is defined to be
$u^{i}(x)\stackrel{{\scriptstyle\rm\Delta}}{{=}}\sum_{\omega\in\Omega}{\phi^{i}_{\omega}(x_{\omega})}.$
The social welfare is defined as the total utility, i.e. $\operatorname{SW}(x)\stackrel{{\scriptstyle\rm\Delta}}{{=}}\sum_{i=1}^{n}{u^{i}(x)}$.
We assume linear project functions and $\theta$-equal sharing unless explicitly said otherwise.
Reconsider the example from Section 1 to illustrate the above model.
Example 1 (Continued).
The scientists $N$ invest in the papers (projects) $\Omega$. Assume that a paper’s total value for the reputation of its authors is proportional to the total investment in the paper. That is, the project’s functions $P_{\omega}$ are linear. In order to be considered an author, a minimum threshold $\theta$ of the maximum contribution is required, and a paper’s total contribution to the authors’ reputation is equally divided between all its authors. This is a shared effort game with a threshold $\theta\in(0,1)$ and equal sharing.
3 Pure Nash Equilibrium
We begin with general results on equilibria in shared effort games. Then, we completely characterise the existence of Nash equilibria for two players, following with existence results for any number of players in Section 3.1. Finally, we provide the prices of anarchy and stability in Section 3.2. We study pure NE in this section, which are natural to play, since a pure shared effort game already has uncountably infinite strategy spaces and non-continuous utilities.
Without a threshold ($\theta=0$), or when $\theta=1$, things would be simple, as we now describe.
Theorem 1.
A game with $M_{\text{eq}}^{0}$ admits a potential function, a pure NE exists and $\operatorname{PoA}=\operatorname{PoS}=1$.
Proof. The strategy space of player $i$ is $S^{i}$, and denote $S\stackrel{{\scriptstyle\rm\Delta}}{{=}}S^{1}\times\ldots\times S^{n}$. Define $P\colon S\to\mathbb{R}$ by $P(x)\stackrel{{\scriptstyle\rm\Delta}}{{=}}\sum_{\omega\in\Omega}{\frac{P_{\omega}(x_{\omega})}{\left|N_{\omega}\right|}}$. This is a potential function, because it is equal to the utility of any player, and therefore, when player $i$ changes her strategy, her utility changes exactly as the potential does.
The game possesses a pure NE, whenever the potential function attains its maximum. As the linear project functions are continuous and the spaces are non-empty compact, being simplexes, so the potential function always attains its maximum (see Lemma $4.3$ in [36]).
Moreover, a profile is an NE if and only if all the players arbitrarily divide their budgets between the most valuable projects. Therefore, being an NE is equivalent to maximizing the social welfare, implying $\operatorname{PoA}=\operatorname{PoS}=1$. ∎
To characterise existence and efficiency of NE in other cases, we need some definitions. Given a strategy profile, we call a project that receives no contribution a vacant project. We define players that do not obtain a share from a given project as dominated at that project. We call them suppressed if they also contribute to that project. Formally,
Definition 2.
The dominated players at a project $\omega$ are $D_{\omega}\stackrel{{\scriptstyle\rm\Delta}}{{=}}N_{\omega}\setminus N^{\theta}_{\omega}$, and the suppressed players at a project $\omega$ are $S_{\omega}\stackrel{{\scriptstyle\rm\Delta}}{{=}}\left\{i\in N_{\omega}:x_{\omega}^{i}>0\right\}\setminus N^{\theta}_{\omega}$. If a player is dominated/suppressed at all the projects where he contributes, we simply call him dominated/suppressed.
In an NE, a player is suppressed at a project if and only if it is suppressed at all the projects where it contributes. This holds since if a player is suppressed at project $p$ but it also contributes to project $q\neq p$ and is not suppressed there, then it would like to move its contribution from $p$ to project $q$.
Consider distinct projects functions $\alpha_{m=m_{1}}>\alpha_{m_{2}}>\alpha_{m_{3}}>\ldots>\alpha_{1=m_{l}}$, denoting $m_{1}=m$ and $m_{l}=1$, where $l\geq 1$ is the number of distinct project functions. We choose $m_{j}$ as an arbitrary $j^{\prime}$ such that $\alpha_{j^{\prime}}=\alpha_{m_{j}}$. For each $j\in\left\{1,2,\ldots,l\right\}$, we call every project $\omega$ with $\alpha_{\omega}=\alpha_{m_{j}}$ a $j$-level project. In the same spirit, consider distinct budgets $B_{n=n_{1}}>B_{n_{2}}>B_{n_{3}}>\ldots>B_{1=n_{p}}$, denoting $n_{1}=n$ and $n_{p}=1$, $p\geq 1$ being the number of distinct budgets. We choose $n_{i}$ as an arbitrary $i^{\prime}$ such that $\alpha_{i^{\prime}}=\alpha_{n_{i}}$. For each $q\in\left\{1,2,\ldots,p\right\}$, call every player $i$ with $B_{i}=B_{n_{q}}$ a $q$-level player, and call $B_{i}$ the $q$-level budget. Thus, given a level, one can talk about projects and budgets on that level.
For each level $j\in\left\{1,2,\ldots,l\right\}$, let $r_{j}$ be the number of projects of level $j$, namely $r_{j}\stackrel{{\scriptstyle\rm\Delta}}{{=}}\left|\left\{\omega\in\Omega:\alpha_{\omega}=\alpha_{m_{j}}\right\}\right|$. Similarly, for each level $j\in\left\{1,2,\ldots,p\right\}$, let $s_{j}$ be the number of players of level $j$, namely $s_{j}\stackrel{{\scriptstyle\rm\Delta}}{{=}}\left|\left\{i\in N:B_{i}=B_{n_{i}}\right\}\right|$. We are now ready to characterise all the equilibria without suppressed players for threshold $\theta=1$.
Theorem 2.
A game with $M_{\text{eq}}^{1}$ possesses NE where nobody is suppressed if and only if all the following conditions hold:
  1. 1.
    The number of distinct project functions $l$ is at least the number of distinct budgets $p$, namely $l\geq p$.
  2. 2.
    For each level $j\in\left\{1,2,\ldots,p-1\right\}$, there either exists a single $j$-level budget such that $B_{n_{j}}>r_{j}B_{n_{j+1}}$, or there exist at least $r_{j}$ budgets of level $j$.
  3. 3.
    For each level $j\in\left\{1,2,\ldots,p-1\right\}$ and level $d\in\left\{j+1,j+2,\ldots,p\right\}$, $\alpha_{m_{j}}\geq(1+\left\lceil{s_{d}/r_{d}}\right\rceil)\alpha_{m_{d}}$.
Proof. In one direction, we assume the existence of $x\in\text{NE}$ where no player is suppressed and prove that all the conditions above hold. The key observation is that since exactly the maximum contributors equally share the project’s value, contributing to a project yields the same value, regardless how many others contribute there too, if nobody is suppressed. Therefore, in an equilibrium without suppression, everyone contributes only to the most profitable project, where that is not suppressed. Thus, all the $1$-level budgets go to the $1$-level projects. The $2$-level budgets would be suppressed if they went where a $1$-level budget is, violating the assumption of not being suppressed, and otherwise such a suppression would be a profitable deviation for any $1$-level player, violating the equilibrium assumption, so they all go to the $2$-level projects. Inductively, $j$-level projects contribute all to $j$-level projects. Therefore, $l\geq p$, yielding condition 1.
Next, consider any level $j\in\left\{1,2,\ldots,p-1\right\}$. If there exists just one $j$-level budget $B_{n_{j}}$, then the only way to render the deviation of any $j+1$-level to a $j$-level project non-profitable is to have more than $B_{n_{j+1}}$ contributed to every $j$-level project, implying $B_{n_{j}}>r_{j}B_{n_{j+1}}$. On the other hand, if there are multiple $j$-level budgets, then no splitting of these budgets is possible, to avoid a suppression from another $j$-level budget being a profitable deviation. Thus, the only way to render the deviation of any $j+1$-level to a $j$-level project non-profitable is to have a full $j$-level budget contribute to every $j$-level project. This implies condition 2.
Finally, consider again any level $j\in\left\{1,2,\ldots,p-1\right\}$ and any greater level $d\in\left\{j+1,j+2,\ldots,p\right\}$. At the $d$ level, at least one project obtains at least $\left\lceil{s_{d}/r_{d}}\right\rceil$ contributions from $d$-level players. Therefore, making the suppression of the total $d$-level contribution to a $d$-level project by a $j$-level player being non-profitable requires $\alpha_{m_{j}}x\geq\alpha_{m_{d}}(x+\left\lceil{s_{d}/r_{d}}\right\rceil B_{n_{d}}),\forall x\in(B_{n_{d}},B_{n_{j}})$, namely $\alpha_{m_{j}}\geq(1+\left\lceil{s_{d}/r_{d}}\right\rceil)\alpha_{m_{d}}$, implying condition 3.
Conversely, we now assume all the conditions above and demonstrate an equilibrium profile without suppressed players. For each level $j\in\left\{1,\ldots,p\right\}$, let all the players of level $j$ contribute all their budgets to the projects of level $j$. There exist enough project levels because of condition 1. If there exists a single such player, she should equally split her budget among all the $j$-level projects. If multiple such players exist, they should not split their respective budgets, and every $j$-level project should receive at least one such player’s contribution, which is possible, because of condition 2. No project of level $d$ should have more than $\left\lceil{s_{d}/r_{d}}\right\rceil$ players contribute there. Intuitively, this means balancedness.
We now prove that the defined profile is an equilibrium without suppression. Suppression does not exist, since a player contributes to the same project only together with equally sized contributions.
As for the non-existence of profitable deviations, consider the possible deviations. Contributing to a more profitable project (a smaller level) would result in suppression, because every project of level $j$ receives a contribution greater than $B_{n_{j+1}}$. Contributing to another project on the same level would either result in suppression if contributing less than the full budget with a full budget contribution, or yield the same utility. Last, contributing to a least profitable (higher level) project is non-profitable because of condition 3, and since no project of level $d$ has more than $\left\lceil{s_{d}/r_{d}}\right\rceil$ players contribute there. This implies the profile is an NE. ∎
Next, we use the structural insights from the previous theorem to analyse the efficiency of equilibria for $\theta=1$.
Proposition 1.
If a game with $M_{\text{eq}}^{1}$ possesses NE where nobody is suppressed, then the social welfare of any such NE is $\sum_{q\leq p}{\alpha_{m_{q}}(\sum_{i:B_{i}=B_{n_{q}}}{B_{i}})}$, whereas the social optimal is $\alpha_{m_{1}}\sum_{i\in N}{B_{i}}$.
Proof. As we showed in the proof of Theorem 2, in any such equilibrium, all the players of level $j$ contribute everything to the projects of the level $j$. This implies the statement. ∎
3.1 Existence of Nash Equilibrium
All Nash equilibria are invariant to multiplication.
Proposition 2.
Consider a game with project functions $P_{j}(x)=\alpha_{j}\cdot x$. If profile $x=\left(x^{1},\ldots,x^{n}\right)$ is an NE, then the following hold:
Multiplying projects:
$x$ is also an equilibrium in the game obtained by multiplying all the project coefficients by some positive $p$.
Multiplying budgets and profile:
$p\cdot x$ is an NE in the game obtained by multiplying all the budgets by some positive $p$.
Proof. The invariance to multiplying the projects stems from the fact that multiplying all the project coefficients by $p$ multiplies all the utilities by $p$. Since this happens to all the utilities, the same relations keep holding between the various strategy profiles, and thus, the same NE remain.
We prove the second part by contradiction. If $p\cdot x$ is not an NE, then there exists a unilateral profitable deviation by player $i$. Denote the profile after such a deviation by $x^{\prime}$. Then, $1/p\cdot x^{\prime}$ is a legal profile in the original game, and it is a unilateral deviation from profile $x$ by player $i$. This deviation is profitable to $i$, since the original deviation is profitable and all the utilities are multiplied by multiplying the profile. This contradicts the assumption that $x$ is an NE of the original game. ∎
Having dealt with $\theta=0$ and $\theta=1$, we may assume $\theta\in(0,1)$. We first assume $2$ players, i.e., $n=2$, and completely characterise this case. We introduce Lemmas 1, 2, and 3, before characterizing the existence of NE. These lemmas describe what must hold in any NE. Their proofs appear in the appendix.
Lemma 1.
Consider an equal $\theta$-sharing game with two players with $10<\theta<$ 1. Then the following hold in any NE.
  1. 1.
    At least one player contributes to a steep project.
  2. 2.
    Suppose that a non-suppressed player, contributing to a steep project, contributes to a non-steep project as well. Then, it contributes either alone or precisely the least amount it should contribute to achieve a portion in the project’s value.
The following lemma treats budgets that are close to each other.
Lemma 2.
Consider an equal $\theta$-sharing game with two players with $10<\theta<$ 1. If $B_{1}\geq\theta B_{2}$, the following hold in any NE.
  1. 1.
    Each player contributes to every steep project.
  2. 2.
    A non-steep project receives the contribution of at most one player.
We need another definition.
Definition 3.
A $2$-steep project is a most valuable non-steep project.
The following lemma treats budgets that are far from each other.
Lemma 3.
Consider an equal $\theta$-sharing game with two players with $10<\theta<$ 1. If $B_{1}<\theta B_{2}$, then the following hold in any NE where no player is suppressed.
  1. 1.
    Player $1$ contributes only to non-steep projects.
  2. 2.
    Each player receives a (strictly) positive utility, unless all projects are the same. [^6] That is, unless $k=m$.
  3. 3.
    Player $2$ contributes alone to every steep project, and perhaps to a non-steep project together with $i$, the threshold amount.
  4. 4.
    The projects that are non-steep and also non-$2$-steep receive zero contribution.
  5. 5.
    If player $2$ contributes to a $2$-steep project, then there exists only a single $2$-steep project.
We are finally ready to characterise the existence of an NE for two players. The characterisation depends on the threshold, the ratio between the budgets and the highest coefficients.
Theorem 3.
Consider an equal $\theta$-sharing game with two players with budgets $B_{1},B_{2}$. W.l.o.g., $B_{2}\geq B_{1}$. Assume $10<\theta<$ 1, and project functions $P_{j}(x)=\alpha_{j}\cdot x$ with coefficients $\alpha_{m}=\alpha_{m-1}=\ldots=\alpha_{m-k+1}>\alpha_{m-k}\geq\alpha_{m-k-1}\geq\ldots\geq\alpha_{1}$ (ordered w.l.o.g.). A pure NE exists if and only if one of the following holds. [^7] If $\alpha_{m-k-1}$ (and) or $\alpha_{m-k}$ does not exist, consider the containing condition to be vacuously true.
  1. 1.
    $B_{1}\geq\theta B_{2}$ and the following both hold.
    1. (a)
      $\frac{1}{2}\alpha_{m}\geq\alpha_{m-k}$,
    2. (b)
      $B_{1}\geq k\theta B_{2}$;
  2. 2.
    $B_{1}<\theta B_{2}$ and also at least one of the following holds.
    1. (a)
      $B_{1}<\frac{\theta B_{2}}{k}$ and $\frac{\alpha_{m-k}}{\alpha_{m}}\leq\min\left\{\frac{1}{1+\theta},\frac{2\theta}{1+\theta}\right\}$,
    2. (b)
      $B_{1}<\frac{\theta B_{2}}{k+\theta^{2}}$ and $\alpha_{m-k}\geq 2\alpha_{m-k-1}$ and $\frac{2\theta}{1+\theta}\leq\frac{\alpha_{m-k}}{\alpha_{m}}\leq\frac{2(1-\theta)}{2-\theta}$ and $m-k$ is the only $2$-steep project,
    3. (c)
      $B_{1}<\frac{\theta}{\left|\Omega\right|}B_{2}$ and all the project functions are equal, i.e. $\alpha_{m}=\alpha_{1}$.
The idea of the proof is as follows. To show existence of an equilibrium under the assumptions of the theorem, we provide a strategy profile and prove that no unilateral deviation is profitable. We show the other direction by assuming that a given profile is an NE and deriving the asserted conditions, employing Lemmas 1, 2 and 3 that describe what holds in an equilibrium.
We prepend the proof of the theorem with the following technical lemma.
Lemma 4.
Consider an equal $\theta$-sharing game with two players with budgets $B_{1},B_{2}$. W.l.o.g., $B_{2}\geq B_{1}$. Assume $10<\theta<$ 1, and project functions $P_{j}(x)=\alpha_{j}\cdot x$ with coefficients $\alpha_{m}=\alpha_{m-1}=\ldots=\alpha_{m-k+1}>\alpha_{m-k}\geq\alpha_{m-k-1}\geq\ldots\geq\alpha_{1}$ (ordered w.l.o.g.). Assume that no player is suppressed anywhere, and player $j$ does not contribute to a non-steep project p. Consider player $i\neq j$.
Then, the following hold.
  1. 1.
    If $\frac{1}{2}\alpha_{m}\geq\alpha_{p}$, then it is not profitable for $i$ to move any budget $\delta>0$ from any subset of the steep projects to $p$ (or to a set of such projects).
  2. 2.
    If $\frac{1}{2}\alpha_{m}>\alpha_{p}$, then it is (strictly) profitable for $i$ to move any budget $\delta>0$ from $p$ to any subset of the steep projects. If $j$ is suppressed after such a move, then requiring $\frac{1}{2}\alpha_{m}\geq\alpha_{p}$ is enough.
  3. 3.
    If $\frac{1}{2}\alpha_{m}<\alpha_{p}$ and it is possible to move $\delta>0$ from any subset of the steep projects to $p$, such that $i$ received and still receives half of the value of these steep projects, then it is (strictly) profitable for $i$.
We are set to prove the theorem.
Proof. ($\Rightarrow$) We prove the existence of NE under the conditions of the theorem. We begin with case 1, supposing that $B_{1}\geq k\theta B_{2}$ and $\frac{1}{2}\alpha_{m-k+1}\geq\alpha_{m-k}$. Let both players allocate $1/k$th of their respective budgets to each of the steep projects. We prove here that this is an NE. This profile provides each player with $k\cdot\frac{1}{2}\alpha_{m}\cdot\frac{B_{1}+B_{2}}{k}=\frac{1}{2}\alpha_{m}\cdot(B_{1}+B_{2})$. For any player $i$, moving $\delta>0$ to some non-steep projects is not profitable, according to part 1 of Lemma 4. Another possible deviation is reallocating budget among the steep projects. Since $B_{1}\geq k\theta B_{2}$, we conclude that $B_{2}\leq\frac{B_{1}}{k\theta}$, so $2$ is not able to suppress $1$ (and the other way around is clearly impossible, even more so) and therefore, merely reallocating among the steep projects will not increase the utility. The only deviation that remains to be considered is simultaneously allocating $\delta>0$ to some non-steep projects and reallocating the rest of the budget among the steep ones. Any such potentially profitable deviation can be looked at as two consecutive deviations: first allocating $\delta>0$ to some non-steep projects, and then reallocating the rest of the budget among the steep ones. Part 1 of Lemma 4 shows that bringing back all $\delta>0$ from non-steep projects to the steep ones, without getting suppressed anywhere (which is possible since $B_{1}\geq\theta B_{2}$) will bear a non-negative profit. Therefore, we can ignore the last form of deviations. Therefore, this is an NE.
We now move to handle case 2. Case 2a: suppose that $B_{1}<\frac{\theta B_{2}}{k}$ and $\frac{\alpha_{m-k}}{\alpha_{m}}\leq\min\left\{\frac{1}{1+\theta},\frac{2\theta}{1+\theta}\right\}$. Let player $1$ invest all its budget in $m-k$, and let $2$ invest $\frac{B_{2}}{k}$ in each steep project. We prove this is an NE. The only possibly profitable deviation for player $1$ is to invest in steep projects. However, since $B_{1}<\frac{\theta B_{2}}{k}$, player $1$ would obtain nothing from the steep projects. Also player $2$ would not gain from a deviation, because first, from our assumption,
$\frac{\alpha_{m-k}}{\alpha_{m}}\leq\frac{1}{\theta+1}\iff\alpha_{m}(B_{1}/\theta)\geq\alpha_{m-k}(B_{1}(1+1/\theta)),$
and therefore, player $2$ would not profit from suppressing player $1$ at project $m-k$. Second, according to our assumption,
$\frac{\alpha_{m-k}}{\alpha_{m}}\leq\frac{2\theta}{\theta+1}\iff\alpha_{m}(\theta B_{1})\geq\frac{\alpha_{m-k}((1+\theta)B_{1})}{2},$
and therefore, player $2$ would not profit from getting a half of the value of project $m-k$. Thus, no deviation is profitable. Therefore this is an NE.
Case 2b: suppose that $B_{1}<\frac{\theta B_{2}}{k+\theta^{2}}$ and $\alpha_{m-k}\geq 2\alpha_{m-k-1}$ and $\frac{2\theta}{1+\theta}\leq\frac{\alpha_{m-k}}{\alpha_{m}}\leq\frac{2(1-\theta)}{2-\theta}$ and $m-k$ is the only $2$-steep project. Let player $1$ invest all its budget in $m-k$, and let $2$ invest $\theta B_{1}$ in $m-k$ and $\frac{B_{2}-\theta B_{1}}{k}$ in each steep project. We prove that this is an NE. The possibly profitable deviations for player $1$ is to invest in steep projects or in project $m-k-1$. Here, we show them to be non profitable. First, since $B_{1}<\frac{\theta B_{2}}{k+\theta^{2}}\iff B_{1}<\theta\frac{B_{2}-\theta B_{1}}{k}$, there is no profit for $1$ from investing in a steep project. Second, according to our assumption,
$\alpha_{m-k}$$\geq$ $2\alpha_{m-k-1}$
$\iff\frac{\alpha_{m-k}(B_{1}(1+\theta))}{2}$$\geq$ $\frac{\alpha_{m-k}(B_{1}(\theta^{2}+\theta))}{2}+\alpha_{m-k-1}(B_{1}(1-\theta^{2})),$
and therefore, player $1$ would not profit from investing in $m-k-1$. Next, we show that also player $2$ does not have incentives to deviate. Since contributing to the non-$2$-steep projects would not increase $2$’s utility, and since the way how the contribution is divided between the steep projects does not influence the utility, the possible deviations to increase player $2$’s utility are transferring budget from the steep projects to $m-k$ or the other way around. We show now that they are not profitable. First, by our assumption,
$\frac{\alpha_{m-k}}{\alpha_{m}}$$\leq$ $\frac{2(1-\theta)}{2-\theta}$
$\iff\alpha_{m}(B_{1}/\theta-\theta B_{1})+\frac{\alpha_{m-k}(B_{1}(1+\theta))}{2}$$\geq$ $\alpha_{m-k}(B_{1}(1+1/\theta)),$
and therefore, player $2$ would not profit from suppressing $1$ on project $m-k$. Second, according to our assumption,
$\frac{\alpha_{m-k}}{\alpha_{m}}\geq\frac{2\theta}{1+\theta}\iff\frac{\alpha_{m-k}(B_{1}(1+\theta))}{2}\geq\alpha_{m}(\theta B_{1}),$
and therefore, player $2$ would not profit from moving $\theta B_{1}$ from $m-k$ to a steep project. Thus, no deviation is profitable. Therefore this is an NE.
Case 2c: suppose that $B_{1}<\frac{\theta}{\left|\Omega\right|}B_{2}$ and all the project functions are equal. Then, player $2$ investing $\frac{B_{2}}{\left|\Omega\right|}$ in every project, and player $1$ using any strategy is an NE. To see this, notice that player $2$ obtains $\alpha_{m}(B_{2}+\sum_{\omega\in\Omega}{x^{1}_{\omega}})$, that is the maximum possible utility. Player $1$ will be suppressed in any attempt to invest, and therefore has no incentive to deviate. Therefore this is an NE.
($\Leftarrow$) We show the other direction now. We assume that a given profile is an NE and derive the conditions of the theorem. We first suppose that $B_{1}\geq\theta B_{2}$ and we shall derive that the conditions of 1 hold.
Since $B_{1}\geq\theta B_{2}$, then according to Lemma 2, each player contributes to every steep project. Suppose to the contrary that $\frac{1}{2}\alpha_{m-k+1}<\alpha_{m-k}$. Let $i$ be a player who contributes to $m$ more than its threshold there, and let $j$ be the other player. Then, by part 3 of Lemma 4, all non-steep projects with coefficients larger than $0.5\alpha_{m}$ must get a positive contribution from $j$, for otherwise $i$ would profit by transferring there part of its budget from $m$. Therefore, the non-steep projects with coefficients larger than $0.5\alpha_{m}$ receive no contribution from $i$, according to Lemma 2.
Therefore, at all the steep projects, player $j$ contributes exactly its threshold value, while $i$ contributes above it. Also, $i$ contributes nothing to any non-steep project: we have shown this for the non-steep projects with coefficients larger than $0.5\alpha_{m}$, now we show it for the rest. If $i$ contributed to a non-steep project with coefficient at most $0.5\alpha_{m}$, he would benefit from deviating to a steep one, by part 2 of Lemma 4 (when the coefficient is exactly $0.5\alpha_{m}$, we use the fact that $j$ would be suppressed by such a deviation).
We assume that $B_{1}\geq\theta B_{2}$, and thus, for any $i\neq j$ we have
$\theta B_{j}\leq B_{i}\iff B_{j}-\theta B_{i}\leq\frac{B_{i}}{\theta}-\theta B_{i}\iff B_{j}-\theta B_{i}\leq\frac{B_{i}-\theta^{2}B_{i}}{\theta}.$
Thus, a non-steep project with coefficients larger than $0.5\alpha_{m}$ receives from $j$ at most $\frac{B_{i}-\theta^{2}B_{i}}{\theta}$, and since $i$ can transfer to that project $B_{i}-\theta^{2}B_{i}$ without losing a share at the steep projects, $i$ can transfer exactly $\theta$-share of $j$’s contribution there and profit thereby, by part 3 of Lemma 4 (that lemma assumes $j$ does not contribute to those non-steep projects, but contributing exactly the threshold to such a project is not worse than alone). This profitable deviation contradicts our assumption and we conclude that $\frac{1}{2}\alpha_{m}\geq\alpha_{m-k}$.
It is left to prove that $B_{1}\geq k\theta B_{2}$. Part 2 of Lemma 4 implies there are no contributions to non-steep projects, since they would render the deviation to the steep projects profitable, unless $\frac{1}{2}\alpha_{m-k+1}=\alpha_{m-k}$, in which case a $2$-steep project can get a positive investment from one player. Thus, the players’ utility is at most the same as when each steep project obtains contributions from both players, and other projects receive nothing. Thus, each player’s utility is at most $k\cdot(\alpha_{m}/2)(\frac{B_{1}+B_{2}}{k})=(\alpha_{m}/2)(B_{1}+B_{2})$. If $2$ could deviate to contribute all $B_{2}$ to a steep project while suppressing $1$ there, player $2$ would obtain $\alpha_{m}(B_{2}+y)$, for some $y>0$. This is always profitable, since
$B_{2}\geq B_{1}\Rightarrow B_{2}+2y>B_{1}\iff\alpha_{m}(B_{2}+y)>(\alpha_{m}/2)(B_{1}+B_{2}).$
Thus, since we are in an NE, $2$ may not be able to suppress $i$ and therefore $B_{2}\leq\frac{B_{1}}{k}\frac{1}{\theta}\Rightarrow B_{1}\geq k\theta B_{2}$. Thus, we have proved that Conditions 1 hold.
Suppose that $B_{1}<\theta B_{2}$ and we shall derive that Conditions 2 hold.
We exhaust all the possibilities for an NE, namely: $1$ is suppressed, $1$ is not suppressed and player $2$ does not contribute to non-steep projects, and $1$ is not suppressed and player $2$ contributes to non-steep projects. We show that each of this options entails at least one of the sub-conditions of 2.
First, assume $1$ is suppressed. [^8] See definition 2. Then, $2$ invests more than $B_{1}/\theta$ at each project. Therefore, $B_{1}<\frac{\theta}{\left|\Omega\right|}B_{2}$. If not all the projects were steep, then $2$ would profitably transfer some amount to a steep project from the non-steep ones, while still dominating $1$ everywhere. This deviation would contradict the profile being an NE. Therefore, all the projects are steep and condition 2c holds.
Assume now that no player is suppressed. Therefore, according to Lemma 3, player $1$ contributes only to the $2$-steep projects, and player $2$ contributes to all the steep ones, and perhaps to a $2$-steep one as well.
First, we assume that player $2$ does not contribute to non-steep projects and show that it entails condition 2a. Next, we assume that $2$ does contribute to non-steep projects and show that this entails condition 2b.
First, assume that player $2$ does not contribute to non-steep projects. Since player $1$ does not prefer to deviate by contributing exactly the threshold to a steep project, $B_{1}<\theta\frac{B_{2}}{k}$ is true. In an NE, player $2$ would not profit from suppressing player $1$ at a $2$-steep project, and therefore
$\alpha_{m}(B_{1}/\theta)\geq\alpha_{m-k}(B_{1}(1+1/\theta))\iff\frac{\alpha_{m-k}}{\alpha_{m}}\leq\frac{1}{\theta+1}.$
In addition, in an NE, player $2$ would not profit from contributing exactly the threshold at a $2$-steep project, and therefore
$\alpha_{m}(\theta B_{1})\geq\frac{\alpha_{m-k}(1+\theta)B_{1}}{2}\iff\frac{\alpha_{m-k}}{\alpha_{m}}\leq\frac{2\theta}{1+\theta}.$
Thus, we have proved that condition 2a holds.
Assume now player $2$ contributes to non-steep projects. By Lemma 3, $m-k$ is the single $2$-steep project where player $1$ contributes all $B_{1}$, while player $2$ contributes $\theta B_{1}$ there, and he splits the rest of his budget between all the steep projects, yielding a positive contribution to each such project.
Assume that a steep project receives $y>0$ (from player $2$, of course). If player $1$ could achieve the threshold $\theta y$, it would deviate, for the following reasons. We have $\alpha_{m}(B_{1}+y)/2>\alpha_{m-k}(B_{1}+\theta B_{1})/2$, unless, perhaps, if $y<\theta B_{1}$. In such a case, however, $1$ can suppress player $2$ and obtain $\alpha_{m}(B_{1}+y)$, which is larger than $\alpha_{m-k}(B_{1}(1+\theta))/2$. Consequently, from the profile being an NE, we conclude that $1$ is not able to achieve the threshold $\theta y$, and therefore
$B_{1}<\theta\frac{B_{2}-\theta B_{1}}{k}\iff B_{1}<\frac{\theta B_{2}}{k+\theta^{2}}.$
In addition, since player $1$ does not prefer to contribute to $m-k$ only the threshold $\theta^{2}B_{1}$ and move the rest to $m-k-1$, it must hold that
$\frac{\alpha_{m-k}(B_{1}(1+\theta))}{2}$$\geq$ $\frac{\alpha_{m-k}(B_{1}(\theta^{2}+\theta))}{2}+\alpha_{m-k-1}(B_{1}(1-\theta^{2}))$
$\iff\alpha_{m-k}$$\geq$ $2\alpha_{m-k-1}.$
Since player $2$ does not want to suppress $1$ at $m-k$, we conclude that
$\alpha_{m}((B_{1})/\theta-\theta B_{1})+\frac{\alpha_{m-k}(B_{1}(1+\theta))}{2}$$\geq$ $\alpha_{m-k}(B_{1}(1+1/\theta))$
$\iff\frac{\alpha_{m-k}}{\alpha_{m}}$$\leq$ $\frac{2(1-\theta)}{2-\theta}.$
Finally, since player $2$ does not prefer moving $\theta B_{1}$ to a steep project over leaving it at $m-k$, it holds that
$\frac{\alpha_{m-k(B_{1}(1+\theta))}}{2}\geq\alpha_{m}(\theta B_{1})\iff\frac{\alpha_{m-k}}{\alpha_{m}}\geq\frac{2\theta}{1+\theta}.$
Therefore, condition 2b holds. To conclude, at least one of the sub-conditions of 2 holds, finalizing the proof of the other direction of the theorem. ∎
Consider some structural insights from the proof of Theorem 3. When the ratio between the budgets is small, namely $B_{1}\geq\theta B_{2}$, both players can be in equilibrium by equally investing in every steep project. When the budgets are further apart, the level-$2$ player has to invest in a level-$2$ project, while the level-$1$ player divides her budget equally between the steep projects. This fits the case of $\theta=1$, analysed in Theorem 2, where the players of each level spread their budgets approximately evenly among the projects of the corresponding level. If the budgets are even further apart and the level-$2$ projects are a bit more profitable than in the former case, then the level-$1$ player can obtain her part of the level-$2$ project, while spreading the rest of her budget evenly among the steep projects. Finally, if the level-$1$ budget allows complete domination, then level-$1$ player can spread her budget evenly, thereby guaranteeing an equilibrium. Thus, there is a tendency to spread budgets in an even manner, though the players with larger budgets often “push” the other players from the more profitable projects. Budgets serve as keys to the more profitable projects.
We conclude that besides the equilibria with $\alpha_{m}=\alpha_{m-k}$, there exists an NE if and only if $\alpha_{m-k}$ is at most a constant fraction of $\alpha_{m}$.
Corollary 1.
Assume the conditions of Theorem 3 and that there exist two projects. Then, once all the parameters besides ${\alpha_{m-k}}$ and ${\alpha_{m}}$ are set and all the conditions involving the budgets are fulfilled, there exists a $C>0$, such that an NE exists if and only if $\frac{\alpha_{m-k}}{\alpha_{m}}\leq C$, and, perhaps, if and only if $\frac{\alpha_{m-k}}{\alpha_{m}}=1$.
Proof. First, for $2$ projects, the condition involving $\alpha_{m-k-1}$ in 2b is vacuously true.
Consider the bounds on the possible values of $\frac{\alpha_{m-k}}{\alpha_{m}}$ such that at least one NE exists, besides case 2c. From Theorem 3, the only way that this corollary could be wrong would require the upper bound on $\frac{\alpha_{m-k}}{\alpha_{m}}$ from 2a to be strictly smaller than the lower bound from 2b, while the two bounds on $\frac{\alpha_{m-k}}{\alpha_{m}}$ from 2b gave a non-empty segment. These conditions mean that both $\frac{1}{1+\theta}<\frac{2\theta}{1+\theta}$ and $\frac{2\theta}{1+\theta}\leq\frac{2(1-\theta)}{2-\theta}$ should hold. The first inequality means $\theta>0.5$, while the second one means $\theta\leq 0.5$. Since these conditions cannot hold simultaneously, the corollary is never wrong. ∎
The sufficiency conditions 1 and 2c extend for a general $n$ as follows.
Theorem 4.
Consider an equal $\theta$-sharing game with $n\geq 2$ players with budgets $B_{n}\geq\ldots\geq B_{2}\geq B_{1}$ (the order is w.l.o.g.), $10<\theta<$ 1, and project functions $P_{j}(x)=\alpha_{j}\cdot x$ with coefficients $\alpha_{m}=\alpha_{m-1}=\ldots=\alpha_{m-k+1}>\alpha_{m-k}\geq\alpha_{m-k-1}\geq\ldots\geq\alpha_{1}$ (the order is w.l.o.g.).
This game has a pure strategy NE if one of the following holds. [^9] If $\alpha_{m-k}$ does not exist, consider the containing condition to be vacuously true.
  1. 1.
    $B_{n-1}\geq\theta B_{n}$ and the following both hold.
    1. (a)
      $\frac{1}{n}\alpha_{m}\geq\alpha_{m-k}$,
    2. (b)
      $B_{1}\geq k\theta B_{n}$;
  2. 2.
    $B_{n-1}<\theta B_{n}$ and also the following holds.
    1. (a)
      $B_{n-1}<\frac{\theta}{\left|\Omega\right|}B_{n}$ and all the project functions are equal, i.e. $\alpha_{m}=\alpha_{1}$.
Proof. It is analogous to the proof for $n=2$, noticing the following. In case 1, everyone equally divides her budget among all the steep projects. In case 2, player $n$ dominates everyone else. ∎
3.2 Efficiency
In order to facilitate decisions, it is important to analyse the efficiency of the various Nash Equilibria. We aim to find the price of anarchy ($\operatorname{PoA}$), which is the ratio of a worst NE’s efficiency to the optimum possible one, and the price of stability ($\operatorname{PoS}$), which is the ratio of a best NE’s efficiency to the optimum possible one. We first completely resolve the two-player case.
Theorem 5.
Consider an equal $\theta$-sharing game with two players with budgets $B_{1},B_{2}$. W.l.o.g., $B_{2}\geq B_{1}$. Assume $10<\theta<$ 1, and project functions $P_{j}(x)=\alpha_{j}\cdot x$ with coefficients $\alpha_{m}=\alpha_{m-1}=\ldots=\alpha_{m-k+1}>\alpha_{m-k}\geq\alpha_{m-k-1}\geq\ldots\geq\alpha_{1}$ (the order is w.l.o.g.). [^10] If $\alpha_{m-k}$ does not exist, consider the containing condition to be vacuously true.
  1. 1.
    Assume that $B_{1}\geq\theta B_{2}$ and the following both hold.
    1. (a)
      $\frac{1}{2}\alpha_{m}\geq\alpha_{m-k}$,
    2. (b)
      $B_{1}\geq k\theta B_{2}$;
    Then, there exists a pure strategy NE and there holds: $\operatorname{PoS}=\operatorname{PoA}=1$.
  2. 2.
    Assume that $B_{1}<\theta B_{2}$ and also at least one of the following holds.
    1. (a)
      $B_{1}<\frac{\theta B_{2}}{k}$ and $\frac{\alpha_{m-k}}{\alpha_{m}}\leq\min\left\{\frac{1}{1+\theta},\frac{2\theta}{1+\theta}\right\}$.
      Then, there exists a pure strategy NE and the following holds. $\operatorname{PoS}=\frac{\alpha_{m}B_{2}+\alpha_{m-k}B_{1}}{\alpha_{m}(B_{1}+B_{2})}$. If the case 2b holds as well, then $\operatorname{PoA}=\frac{\alpha_{m}(B_{2}-\theta B_{1})+\alpha_{m-k}(B_{1}(1+\theta))}{\alpha_{m}(B_{1}+B_{2})}$; otherwise, $\operatorname{PoA}=\operatorname{PoS}$.
    2. (b)
      $B_{1}<\frac{\theta B_{2}}{k+\theta^{2}}$ and $\alpha_{m-k}\geq 2\alpha_{m-k-1}$ and $\frac{2\theta}{1+\theta}\leq\frac{\alpha_{m-k}}{\alpha_{m}}\leq\frac{2(1-\theta)}{2-\theta}$ and $m-k$ is the only $2$-steep project.
      Then, there exists a pure strategy NE and the following holds. If the case 2a holds as well, then $\operatorname{PoS}=\frac{\alpha_{m}B_{2}+\alpha_{m-k}B_{1}}{\alpha_{m}(B_{1}+B_{2})}$; otherwise, $\operatorname{PoS}=\frac{\alpha_{m}(B_{2}-\theta B_{1})+\alpha_{m-k}(B_{1}(1+\theta))}{\alpha_{m}(B_{1}+B_{2})}$. In any case, $\operatorname{PoA}=\frac{\alpha_{m}(B_{2}-\theta B_{1})+\alpha_{m-k}(B_{1}(1+\theta))}{\alpha_{m}(B_{1}+B_{2})}$.
    3. (c)
      $B_{1}<\frac{\theta}{\left|\Omega\right|}B_{2}$ and all the project functions are equal, i.e. $\alpha_{m}=\alpha_{1}$.
      Then, there exist pure NE and $\operatorname{PoS}=1,\operatorname{PoA}=\frac{B_{2}}{B_{1}+B_{2}}$.
We derive the exact lower bound (infimum) and the maximum of the price of anarchy and stability.
Corollary 2.
Consider an equal $\theta$-sharing game with two players with budgets $B_{1},B_{2}$. W.l.o.g., $B_{2}\geq B_{1}$. Assume $10<\theta<$ 1, and project functions $P_{j}(x)=\alpha_{j}\cdot x$ with coefficients $\alpha_{m}=\alpha_{m-1}=\ldots=\alpha_{m-k+1}>\alpha_{m-k}\geq\alpha_{m-k-1}\geq\ldots\geq\alpha_{1}$ (the order is w.l.o.g.). [^11] If $\alpha_{m-k}$ does not exist, consider the containing condition to be vacuously true. Then, the infimum of $\operatorname{PoS}$ over all the cases is $\frac{k}{k+\theta}(>0.5)$, and the maximum is $1$. The same holds for $\operatorname{PoA}$.
We next prove some efficiency results for a general $n\geq 2$.
Theorem 6.
Consider an equal $\theta$-sharing game with $n\geq 2$ players with budgets $B_{n}\geq\ldots\geq B_{2}\geq B_{1}$, $10<\theta<$ 1 (the order is w.l.o.g.), and project functions $P_{j}(x)=\alpha_{j}\cdot x$ with coefficients $\alpha_{m}=\alpha_{m-1}=\ldots=\alpha_{m-k+1}>\alpha_{m-k}\geq\alpha_{m-k-1}\geq\ldots\geq\alpha_{1}$ (the order is w.l.o.g.). [^12] If $\alpha_{m-k}$ does not exist, consider the containing condition to be vacuously true.
  1. 1.
    Assume that $B_{n-1}\geq\theta B_{n}$ and the following both hold.
    1. (a)
      $\frac{1}{n}\alpha_{m}\geq\alpha_{m-k}$,
    2. (b)
      $B_{1}\geq k\theta B_{n}$;
    Then, there exist pure NE and $\operatorname{PoS}=1$, $\operatorname{PoA}\geq\frac{(1+(n-1)\theta)(B_{n-1}+B_{n})}{n(\sum_{i\in\left\{1,2,\ldots,n\right\}}{B_{i}})}$.
  2. 2.
    Assume that $B_{n-1}<\theta B_{n}$. Then, $\operatorname{PoA}\geq\frac{B_{n}}{\sum_{i\in\left\{1,2,\ldots,n\right\}}{B_{i}}}$.
    If we also have $B_{n-1}<\frac{\theta}{\left|\Omega\right|}B_{n}$ and all the project functions are equal, i.e. $\alpha_{m}=\alpha_{1}$, then there exist pure NE and $\operatorname{PoS}=1,\operatorname{PoA}=\frac{B_{n}}{\sum_{i\in\left\{1,2,\ldots,n\right\}}{B_{i}}}$.
To prove more bounds on the price of anarchy, we employ the so-called smoothness argument from Roughgarden [46]. This means showing that for a socially optimal profile $x^{*}$ and any profile $x$, there exist $\lambda>0$ and $\mu>-1$ such that
$\sum_{i=1}^{n}{u^{i}((x^{*})^{i},x^{-i})}+\mu\cdot\operatorname{SW}(x)\geq\lambda\cdot\operatorname{SW}(x^{*}).$ (1)
Roughgarden proves that this implies that each pure NE has a social welfare of at least $\frac{\lambda}{1+\mu}$ of the optimal social welfare, i.e. $\operatorname{PoA}\geq\frac{\lambda}{1+\mu}$.
Remark 1.
We employ the relaxation from [46, Remark $2.3$] that $x^{*}$ does not have to be any profile but we may pick any socially optimal one, and we use the payoff-maximization version of the argument (see [46, Section $2.3.2$]).
Equipped with this tool, we prove the following.
Theorem 7.
Consider an equal $\theta$-sharing game with $n\geq 2$ players with budgets $B_{n}\geq\ldots\geq B_{2}\geq B_{1}>0$ (the order is w.l.o.g.), $0\leq\theta\leq 1$, and project functions $P_{j}(x)=\alpha_{j}\cdot x$ with coefficients $\alpha_{m}=\alpha_{m-1}=\ldots=\alpha_{m-k+1}>\alpha_{m-k}\geq\alpha_{m-k-1}\geq\ldots\geq\alpha_{1}>0$ (the order is w.l.o.g.). Let $l\in\left\{1,2,\ldots,n\right\}$ be the smallest integer such that $B_{1}\leq B_{2}\leq\ldots\leq B_{l-1}<\theta B_{n}$, but $B_{l}\geq\theta B_{n}$. Then,
$\operatorname{PoA}\geq\frac{1+(n-1)\theta}{n}\frac{\sum_{i=l}^{n-1}{B_{i}}}{\sum_{i=1}^{n}{B_{i}}}+\frac{1+(n-l)\theta}{n-l+1}\frac{B_{n}}{\sum_{i=1}^{n}{B_{i}}}.$
Proof. We show the smoothness argument for some $\lambda>0$ and $\mu>-1$. Setting $\mu$ to zero implies that the left hand side of Eq. (1) is minimised as follows. The utility of players less than $l$ is zero, since $n$ can suppress them. To minimise the utility of a player $i$ that $l\leq i\leq n-1$, every other player contributes exactly $\theta B_{i}$ (if her budget allows), and to minimise the utility of $B_{n}$, every player from $l$ till $n-1$ contributes $\theta B_{n}$, while the players less than $l$ contribute nothing. This yields the following lower bound on the left hand side:
$\sum_{i=l}^{n-1}{\frac{\alpha_{m}B_{i}(1+(n-1)\theta)}{n}}+\frac{\alpha_{m}B_{n}(1+(n-l)\theta)}{n-l+1}+0\cdot\operatorname{SW}(x).$
The right hand side of Eq. (1), equal to $\lambda\cdot\alpha_{m}\sum_{i=1}^{n}{B_{i}}$, is bounded by the left hand side if and only if
$\lambda\leq\frac{1+(n-1)\theta}{n}\frac{\sum_{i=l}^{n-1}{B_{i}}}{\sum_{i=1}^{n}{B_{i}}}+\frac{1+(n-l)\theta}{n-l+1}\frac{B_{n}}{\sum_{i=1}^{n}{B_{i}}}.$
Combined with $\mu=0$, this implies the lower bound on the price of anarchy. ∎
If $B_{n-1}<\theta B_{n}$, this theorem with $l=n$ implies that the price of anarchy is at least $\frac{B_{n}}{\sum_{i=1}^{n}{B_{i}}}$, matching part 2 of Theorem 6.
To simplify the theorem, we draw a simpler, though a looser, bound.
Corollary 3.
Under the conditions of Theorem 7, the following bound holds. Let $l$ be the first integer such that $B_{1}\leq B_{2}\leq\ldots\leq B_{l-1}<\theta B_{n}$, but $B_{l}\geq\theta B_{n}$. Then $\operatorname{PoA}\geq\theta\frac{\sum_{i=l}^{n}{B_{i}}}{\sum_{i=1}^{n}{B_{i}}}.$
To summarise, we first characterise the cases of zero threshold and threshold $1$, next characterising the existence and efficiency of equilibria for $2$ players and provide some existence and efficiency results for $n\geq 2$ players. Notice that our (pure) model is not the mixing of the model where a player may invest in at most one project, because the mixing would extend the utilities linearly in the mixing coefficients [40, Definition 32.1], which is not the case in our (pure) model with a positive threshold.
4 Narrowing down Nash to Cyclically Strong Equilibrium
We now introduce a new generic method to study solution concepts. Having introduced a new solution concept by relaxing or strengthening some properties of an existing solution concept, imagine it subsequently turns out that both concepts coincide on a subclass of games. Now, if the new concept generally strengthens the previous one, then we conclude that on the subclass where both concepts are equivalent, the requirements of the original concept imply the new concept, implying a narrowing down of the original concept to the new one. On the other hand, if the new concept relaxes the original one, then their equivalence on a certain subset of games means that already the lighter requirements imply the original concept, demonstrating a broadening of the original concept. This also allows characterising one solution using the equivalent one, and to transfer the efficiency bounds.
Here, we are going to (apparently) strengthen the Nash equilibria on thresholded shared effort games by adding the requirement that no cyclical deviation of whole budgets is allowed. Cyclical moves are practically and theoretically important, such as the famous top trading cycles mechanisms [1]. Subsequently, we will prove that the resulting solution concept is actually equivalent to Nash equilibrium on thresholded shared effort games where no budget can suppress another one, thereby constituting a narrowing down of Nash equilibrium. In other words, having no profitable unilateral deviations already implies possessing no cyclical deviation of whole budgets. This exemplifies a new useful technique we hope will serve many researchers in the future.
Formally, given a shared effort game, we define
Definition 4.
A cyclically strong equilibrium is a Nash equilibrium $x$, where there exists no coalitional deviation of players $i_{1},i_{2},\ldots,i_{p}$ such that
  1. 1.
    for each $l=1,2,\ldots,p$, there exists a project $\omega\in\Omega$, such that $x^{i_{l}}_{\omega}=B_{i_{l}}$, namely, all $i_{l}$’s budget is in one project;
  2. 2.
    if each player $i_{l}$ above deviates to invest all $B_{i_{l}}$ into the cyclically next player’s project, thereby creating profile $x^{\prime}$, then no deviating player loses and at least one of them strictly benefits w.r.t. $x$.
Cyclically strong equilibrium is a restriction of Nash and a relaxation of strong Nash equilibrium [6].
Interestingly, the new concept often coincides with Nash equilibrium.
Theorem 8.
In any thresholded shared effort game where $B_{i}\geq\theta B_{j},\forall i,j\in N$, any Nash equilibrium is also cyclically strong.
Proof. Consider an $x\in\text{NE}$, and consider a cyclic coalitional deviation by players $i_{1},i_{2},\ldots,i_{p}$ from Definition 4, assuming by contradiction that no player loses and at least one strictly benefits. Since no unilateral deviation is profitable, while now no player loses and at least one strictly benefits, then either some $B_{i_{l}}<\theta B_{i_{l+1}}$ (modulo $p$), which is assumed not to be the case, or each deviation on this cycle has to be of a budget of size at most the size of the previous deviation, thus $B_{i_{1}}\geq B_{i_{2}}\geq\ldots\geq B_{i_{p}}\geq B_{i_{1}}$, with at least one inequality being strict. However, this cycle implies $B_{i_{1}}>B_{i_{1}}$, in contradiction to the assumption of a cyclic deviation above. ∎
This property allows for a better theoretical study of Nash equilibria, as well as providing a practical property they all possess.
5 Mixed Nash Equilibrium
As we have seen, a game may not possess a pure NE. Therefore, we naturally turn to mixed extensions [^13] A mixed extension has strategies that are distributions on the pure strategies and the respective utilities are the expected utilities under these distributions. and ask whether a mixed extension always has an NE. At first, this is unclear. As the game is infinite, the theorem by Nash [38] about the existence of a mixed NE in finite games is irrelevant. Since the game is not continuous, even the theorem by Glicksberg [18] about the existence of a mixed NE in continuous games is not applicable. Fortunately, we answer affirmatively employing a more general existence theorem by Maskin and Dasgupta [32, Theorem $5^{*}$], which requires some definitions. Let the strategy sets be $A_{i}\subset\mathbb{R}^{m}$. For each pair of players $i,j\in{1,\ldots,n}$, let $D(i)$ be a positive natural, and for a $d\in\left\{1,\ldots,D(i)\right\}$, let $f^{d}_{i,j}\colon\mathbb{R}\to\mathbb{R}$ be one-to-one and continuous, such that $(f^{d}_{i,j})^{-1}=f^{d}_{j,i}$. For every player $i$, we define
$A^{*}(i)\stackrel{{\scriptstyle\rm\Delta}}{{=}}\{(a_{1},\ldots,a_{n})\in A\mid\exists j\neq i,\exists k\in\left\{1,\ldots,m\right\},\exists d\in\left\{1,\ldots,D(i)\right\},$
[…]
They define weakly lower semi-continuity, which intuitively means that there is a set of directions, such that approaching a point from any of these directions gives values at least equal to the function at the point. Find the precise definition in appendix Section A.
Theorem $5^{*}$ from [32] states that if the strategy sets $A_{i}\subset\mathbb{R}^{m}$ are non-empty, convex, and compact, the utility function $u^{i}$ are continuous, except for $A^{**}(i)\subseteq A^{*}(i)$, the sum of all the utilities is upper semi-continuous, and for every player $i$, $u^{i}(a_{i},a_{-i})$ is bounded and weakly lower semi-continuous in $a_{i}$, then, there exists a mixed Nash equilibrium in this game.
Using Theorem $5^{*}$ from [32], we prove (see the proof in Section A)
Theorem 9.
Any (linear) shared effort game with $\theta$-equal sharing has a mixed Nash equilibrium.
The existence result automatically extends to the solution concepts that include mixed Nash equilibria, such as correlated [7] and coarse correlated [37] equilibria. Luckily, not only existence results but also some efficiency bounds extend to other equilibria, as we describe next.
First, the invariance to multiplication holds for the mixed, correlated and coarse correlated equilibria as well, since the proof Proposition 2 extends from linearity of expectation.
We are about to show that some bounds on the social welfare of solution concepts extend to the other equilibria as well. An important preliminary observation is that the maximum social welfare stays the same even when (correlated) randomization is allowed; it is always $\alpha_{m}\sum_{i=1}^{n}{B_{i}}$.
Consider the results of Theorem 6. Its lower bounds on the price of anarchy stem from the utility that certain players can always achieve, and the bounds, therefore, hold for mixed, correlated and coarse correlated equilibria as well. Since any pure NE is also a mixed/correlated/coarse-correlated NE, the rest of the efficiency results, based on presenting an NE, also extend to the other solution concepts.
As for Theorem 7 and its corollary, Roughgarden [46, Theorem $3.2$] proves that the lower bounds on the price of anarchy that are proven using the smoothness argument go over to the mixed, correlated and coarse correlated equilibria.
To conclude, NE exist in the mixed case, and the efficiency bounds from Theorems 6 and 7 apply there too.
6 Related Work
Since understanding motivation is necessary to implement recommendations about contribution, we first discuss why people contribute to projects and how such contributions have been modeled. Then, we provide the basic game-theoretic background and present the existing work on existence and efficiency of NE for sharing effort, concluding that no analysis of the general setting has taken place, a gap which we partially fill.
Contribution to Projects
Motivation to contribute to public projects can be both extrinsic, like a payment or a record for the CV, and intrinsic, such as exercising one’s favorite skills [23] or, mostly, conservation citizen science projects [33]. Allahbakhsh et al. [3] discuss rewards for contributing to crowdsourcing, based on quality, where contributors have to fulfill given requirements and build reputation (extrinsic) over time. Wang et al. [51] model motivation to contribute to online travelling communities and conclude the importance of both the practical motives, such as supporting travellers, building relationships, and hoping for a future repay (extrinsic), as well as of internal drives to participate. Forte and Bruckman [16] study why people contribute to Wikipedia, by asking contributors, and conclude that the reasons are similar to those of scientists and include the desire to publish facts about the world (intrinsic). Bagnoli and Mckee [10] empirically check when people contribute to a public good, like building a playground. They find that if people know the threshold for the project’s success and benefit from collective contributing, then they will contribute, in agreement with the theory of [9]. The work argues that knowing such information is realistic, suggesting the real cases of hiring a lobbyist and paying to a ski club as evidence. This conclusion supports our rationality assumption.
The concrete ways to motivate such contributions have been studied too. For instance, Harper et al. [21] find that explicitly comparing a person’s contribution to the contribution of others helps focusing on the desired features of the system, but does not change the interest in the system per se. The influence of revealing how much people contribute to a movie rating community is experimentally studied in [13]. Initiating participation in online communities is experimentally studied in [29] on the example of the influence of similarity and uniqueness of ratings on participation.
We now present the required theoretical background.
Background: Equilibria and Efficiency
Given a general non-cooperative game in strategic form $(N,(S_{i})_{i\in N},(u_{i})_{i\in N})$, a Nash equilibrium is a profile $s\in S$, such that no unilateral deviation is beneficial, namely $\forall i\in N,\forall s_{i}^{\prime}\in S_{i}$, $u_{i}(s_{i}^{\prime},s_{-i})\leq u_{i}(s)$.
An NE can be inefficient, such as in the famous example of the prisoner’s dilemma [40, Example $16.2$]. The ratios of the objectives in the least or the most efficient NE and in the optimum, called price of anarchy ($\operatorname{PoA}$[26] and price of stability ($\operatorname{PoS}$[4], respectively, constitute the most popular approaches to quantifying this inefficiency [39, Chapter 17]. The price of anarchy measures the best guarantee on an NE, while the price of stability measures the cost of leading the game to a specific equilibrium. Roughgarden and Tardos [39, Chapter 17] discuss inefficiency of equilibria in non-cooperative games and consider the examples of network, load balancing and resource allocation games. This work argues that understanding exactly when selfish behaviour is socially profitable is important, since in many applications, implementing control is extremely difficult.
We combinatorially study existence and efficiency of NE, and then turn to simulations.
Background: Classical Fictitious Play
In our simulations, we generalise and employ the widely studied fictitious play, introduced by Brown [12], to find equilibria. In this play, each player best-responds to the product of cumulative marginal histories of the others’ actions at every time step. It is a myopic learning process. If the game is finite, then if a fictitious play converges, then the distribution in its limit is an NE [30]. Conversely, a game is said to possess the fictitious play property if every fictitious play approaches equilibrium in this game [35]. Many researchers show games that possess this property, for example, finite two-player zero-sum games [45] and finite weighted potential games [35]. A famous example for a game without such property is a $3\times 3$ game by Shapley [48]. In this game, there is a cyclic fictitious play that plays each strategy profile for at least an exponentially growing number of times, and therefore, does not converge at all. Moreover, even its subsequences do not converge to an NE.
Background: Mixed Equilibria and their Existence
Since pure equilibria may not exist, we also consider the mixed extensions of a game, where the strategies are the probability distributions over the original (pure) strategies and the utilities are defined as the expected utilities under the played distributions. Then, a mixed Nash equilibrium is defined as an equilibrium of the mixed extension of a game. Regarding the existence of a mixed NE, Nash proved in his classical paper [38] that a mixed extension of a finite game always possesses an NE. Glikcsberg [18] showed the existence of a mixed NE for continuous games. However, in shared effort games with a positive threshold, the threshold creates discontinuity. Dasgupta and Maskin prove the existence of a mixed equilibrium for a subclass of possibly discontinuous games [32]. We show that shared effort games can be cast to fit Dasgupta and Maskin in Theorem 9.
Extant Models of Sharing Effort
The models that resemble ours, but expressed in the terms of cooperative games, where every coalition of players has a value, include the works by Zick, Elkind and Chalkiadakis [53, 52]. Contributing to a coalition can be considered as contributing to a project in a shared effort game, and in both cases players have budgets and obtain shares. Another coalitional model, where participating in a coalition can be seen as contributing to projects appears in [17], though in addition to being a different kind of game, this model allows no participation in multiple coalitions. Despite these similarities, cooperative games are not games in our sense, since they do not define utilities. Moreover, even if we considered a shared effort game as a particular case of a cooperative game, a positive threshold would vitiate the individual rationality of the core, since a player who obtains a positive share when she is the only contributor to a project may obtain nothing when others contribute to the same project.
We now move to non-cooperative games modelling sharing effort. Shared effort games where only the biggest contributor obtains the project’s value, while everyone pays, are called all-pay auctions, and their equilibria are studied, for instance, by Baye, Kovenock and de Vries [34]. That work shows cases where each player obtains the expected payoff of zero, and cases where the winner obtains the difference between the two highest valuations, while the rest obtains zero. All-pay auctions model lobbying, single-winner contests, political campaigns, striving for a job promotion (see e.g. [49]) and Colonel Blotto games with two [44] or more [28, 11] players. In a Colonel Blotto game, the colonels divide their armies between battlefields, and at every battlefield, the larger force wins. The number of the won battlefields defines the utility of a colonel. Namely, the winner of a battlefield takes it fully to her utility. For two players, Roberson [44] analyses the equilibria of this game and their expected payoffs. Any outcome is socially optimum, since this is a constant-sum game, modelling a confrontation. The $n$-player Blotto game has been studied too [28]. Shared effort games can model these games, but typically, shared effort models cooperation, where a project’s value increases with the total contribution.
The models below feature no thresholds required to share a project’s value. Anshelevich and Hoefer [5] consider an undirected graph model, where the nodes are the players and each player divides its budget between its adjacent edges in minimum effort games (where the edges are $2$-player projects), each of which equally rewards both sides by the value of the project’s success (i.e., duplication instead of division). Anshelevich and Hoefer prove the existence of equilibria, find the complexity of finding an NE, and find that the $\operatorname{PoA}$ is at most $2$. A related setting of multi-party computation games appears in Smorodinsky and Tennenholtz [50]. There, the players are computing a common function that requires them to compute a costly private value, motivating free-riding. The work suggests a mechanism, where honest computation is an NE. This differs from our work, since they consider cost minimization, and the choice of the players is either honestly computing or free riding, without choosing projects. For shared effort games with specific conditions (obtaining at least a constant share of one’s marginal contribution to the project’s value and no contribution threshold, i.e. $\theta=0$), Bachrach et al. [8] show that the price of anarchy ($\operatorname{PoA}$) is at most the number of players. That work also upper bounds the $\operatorname{PoA}$ for the case of convex project functions, where each player receives at least a constant share of its marginal contribution to the project’s value. The academic game [2] considers discrete weights, limits the collaboration to at most two agents per project, and the utility functions model synergy and various ordering of the authors, which effect is studied. This model employs the pairwise stability approach [22], which in addition to guarding against unilateral deviations, guards against $2$ agents deviating to a common project too. Gollapudi et al. [19] allow an agent merely to choose which single team to join. They consider several profit sharing models, including equal sharing, but without any threshold.
There has been no research of the NE of our problem with a $\theta\in[0,1]$ sharing mechanism before the preliminary version [43] of our paper. A positive $\theta\in(0,1]$ contradicts the condition of receiving a constant share. We provide precise conditions for existence of NE, and find their efficiency. There exist two works that variate this model in various directions. In [42], quotas and other requirements from projects, such as paper and grants applications, are modeled by requiring a project to be good enough in order to actually obtain its value. They find equilibria and their efficiency and compare the efficiency of equilibria for being within a quota and for attaining at least a certain minimum value. In [41], investing in reciprocal interactions, like attending seminars and meeting friends, is modeled be assuming each project to be a reciprocal interaction from which the interacting parties obtain a value. They prove the convergence of such a process and analyse the existence and efficiency of the equilibria.
7 Conclusions and Future Work
This paper considers shared effort games where the players contribute to the given projects, and subsequently share the linear values of these projects, conditionally on the allocated effort. We study existence and efficiency of the NE. We discover that multiplying all the project functions or budgets by the same factor does not change the equilibria. We first treat the equilibria for thresholds of 00 and $1$. For threshold 00, pure equilibria always exist and are socially optimal, requiring no regulation. For threshold of $1$, on the other hand, there exist non-trivial conditions for existence of equilibria without suppression, and when they hold, the efficiency is uniquely determined and depends on the ratio between the sums of the budgets weighted by the project functions. Thus, striving for high efficiency would require making the projects similar to one another, but otherwise, regulation will not help, as the efficiency is the same across all such equilibria. We then characterise the existence and efficiency of pure NE for shared effort games with two players. When an NE exists and the budgets are close to each other, all the NE are socially optimal. When the budgets are further apart, in the sense that smallest budget is less than threshold times the largest one, the efficiency depends on the ratio of the budgets and of the two or three largest projects’ coefficients but is always greater than half of the optimum.
When the budgets are close, we demonstrate an optimal NE where everyone equally spreads her budgets between the most valuable projects. This motivates the organisers of any project to make their project most valuable possible. Even second best project can receive no contributions in an optimal NE.
For arbitrarily many players, we find socially optimal pure equilibria in some cases and bound the efficiency from below. We obtain further results on the existence and efficiency of NE for more than two players by simulating fictitious play. Still, the main contribution of the simulations resides in the presented methodology, generalising fictitious play and computing a best response. To this end, we generalise fictitious play to infinite strategy spaces and describe some of the best responses of a player to the other players’ strategies. First, we corroborate all the theoretical predictions for the simulated cases. The most important factor for existence and efficiency of an equilibrium is the ratio of the largest to the second largest project function coefficients and of the largest to the second largest budgets. Therefore, to influence the projects and the players to some extent, one should influence the projects with the highest values and the players with the largest budgets.
Some efficiency bounds persist in the mixed extension, where we show that an NE always exists.
Our main conceptual contribution is narrowing down a Nash equilibrium to a new solution concept, discovering that without suppression of whole budgets, Nash equilibria have no profitable cyclical deviations. That inspires thinking about other interactions and relevant solution concepts, which may coincide on important classes of games, motivating much future work.
Further directions for future work follow. Since the real value usually depends on the total contribution linearly up to a point, we should study piecewise project functions. Sometimes the value of the project grows in discrete steps; for example, in the times required to complete a product or to complete a certain part of an article. In those cases, replacing our linear model with a discrete value function would require us to check which properties still hold, and this more precise model would also inspire further realistic models in game theory. More generally, we would like to extend our complete theoretical characterization of the existence of (pure) NE to more than two players and to various non-linear project functions. Next, we can try finding various NE that are not yet analysed analytically by approximately finding a best response and extending the simulations to more than two projects. Our simulations possess extra range representativeness because of the multiplication invariance, and we have tried to consider the practically realistic ranges parameters. Still, the parameter range of the simulations should be extended based on the studied scenario. Another numerical approach is discretizing the game and applying off-the-shelf software to find the equilibria of the discretization, such as Gambit  [^14] http://gambitproject.readthedocs.io/en/v15.1.1/intro.html# or Game Theory Explorer  [^15] http://banach.lse.ac.uk/. Since randomization can be deliberately undertaken or describe beliefs or uncertain behaviour [40, Section 3.2], we would like to find concrete mixed equilibria to be able to advise on playing them or at least predict the outcomes, like we do here for the pure equilibria. Finally, we would like to vary the model, studying a relative threshold that depends on the median of the contributions, to dampen the influence of extreme contributors, or studying an absolute threshold, and imposing further constraints on the players. Converting important interactions, like investment policies of political campaigns, to our model opens a promising avenue, too.
To conclude, we have analysed when contributions to public projects are in equilibrium and what is lost in the equilibria relatively to the best possible contribution profiles. The theoretical analysis of efficiency implies that for two players with close budgets, no coordination is needed, since the price of anarchy is $1$. The price of anarchy is close to $1$ also for two players with budgets that are far from each other, and it is always more than half. For three or more players, some coordination may substantially improve the total utility, though we have seen many cases with efficiency above $0.75$. We have also provided conditions for a general number of players where every equilibrium is almost optimal, so no coordination is required. In the scenarios where much is lost relatively to the optimum, coordination may improve efficiency.
Appendix A Proofs
We now prove Lemma 1.
Proof. First, at least one player contributes somewhere, since otherwise any positive contribution would be a profitable deviation for every player (all $\alpha_{i}$s and $B_{j}$s are positive). Moreover, at least one of the players contributes to a steep project, for the following reasons. If only the non-steep projects receive a contribution, then take any such project $p$. If a single player contributes there, then this player would benefit from moving to contribute to a vacant steep project. If both players contribute to $p$, then if one is suppressed, it would like to deviate to any steep project where it would not be suppressed, and if no-one is suppressed, then a player who contributes not less would like to contribute to a vacant steep project instead.
We prove part 2 now. Let $i\in N$ be any non-suppressed player among those who contribute to a steep project, w.l.o.g., to project $m$. Assume first that player $j\neq i$ is not suppressed. Then, for any non-steep project where $i$ contributes, $i$ contributes either alone or precisely the least amount it should contribute to achieve a portion in the project’s value, because otherwise $i$ would like to increase its contribution to $m$ on the expense of decreasing its contribution to the considered non-steep project.
Now, consider the case where $j$ is suppressed. Then, even if $j$ contributes to a non-steep project where $i$ contributes (and is suppressed there), $i$ still will prefer to move some budget from this project to $m$, since $i$ receives the whole value of $m$ as well. Thus, this cannot be an NE.
The proof of Lemma 2 appears now.
Proof. Since $B_{1}\geq\theta B_{2}$, no player is suppressed, because any player prefers not being suppressed, and at any project, a player who concentrates all its value there is not suppressed.
Every steep project receives a positive contribution from each player, for the following reasons. If only a single player contributes to a steep project, then the player who does not contribute there will profit from contributing there exactly the threshold value, while leaving at least the threshold values at all the projects where it contributed. There is always a sufficient surplus to reach the threshold because $B_{1}\geq\theta B_{2}$. If no player contributes to a steep project $p$, then there exists another steep project $q$, where two players contribute, according to part 1 of Lemma 1 and what we have just described. The player who contributes there strictly more than the threshold would profit from moving some part of his contribution from $q$ to $p$, still remaining not less than the threshold on $q$, contradictory to having an NE.
We next prove the second part of the lemma. Since both players are non-suppressed contributors to steep projects, then, according to part 2 in Lemma 1, we conclude that there exist no non-steep projects where $j$ and $i$ contribute together. ∎
We now present the proof of Lemma 3.
Proof. We prove part 1 first. Consider an NE profile. Assume to the contrary that player $1$ contributes to a steep project, w.l.o.g., to project $m$. Since $B_{1}<\theta B_{2}$ and no player is suppressed, player $2$ could transfer to $m$ budget from other projects, such that at each project, where $2$ was obtaining a share of the value, $2$ still obtains a share, and $2$ suppresses $1$ at $m$. This would increase $2$’s utility, contrary to the assumption of an NE.
Now, we prove parts 2 and 3. Since $1$ does not contribute to steep projects, part 1 of Lemma 1 implies that $2$ contributes to a steep project (say, $2$ contributes $y>0$ to project $m$), and part 2 of Lemma 1 implies that if $2$ contributes to a non-steep project, it contributes there either alone or precisely the least amount it should contribute to achieve a portion in the project’s value. Since contributing alone is strictly worse than contributing this budget to a steep project, $2$ may only contribute together with $1$, the threshold amount. Therefore, $1$ receives a positive value in this profile, and we have part 2. The only thing left to prove here is that $2$ contributes to each steep project. If not, $1$ would prefer to move some of its contribution there, in contradiction to the assumption of an NE.
We prove part 4 now. Assume to the contrary that a non steep and non $2$-steep project receives a contribution. We proved in part 3 that $1$ contributes there, alone or not. For her, moving a small enough utility to a $2$-steep project would increase her utility, regardless whether $2$ contributes to any of those projects. This is so because if $2$ contributes together with $1$, it contributes precisely the threshold amount, according to part 3. This incentive to deviate contradicts the assumption of an NE.
We prove part 5 now. If $2$ contributes to a $2$-steep project $p$, then there may not exist another $2$-steep project, since otherwise $1$ would like to transfer a small amount from $p$ to another $2$-steep project $q$, such that without losing a share of the value of $p$, player $1$ gets the whole value of project $q$.
The proof of Lemma 4 appears now.
Proof. Before moving, player $i$’s utility is $\sum_{q\in\Omega}{(\frac{1}{2}\text{or }1)\alpha_{q}\cdot\left(x^{1}_{q}+x^{2}_{q}\right)}$.
For part 1, assume $\frac{1}{2}\alpha_{m}\geq\alpha_{p}$. If $i$ moves $\delta>0$ from the steep projects to $p$, then its utility from the steep projects decreases by at least $0.5\alpha_{m}\delta$, and its utility from $p$ increases by $\alpha_{p}\delta$. The total change is $(-0.5\alpha_{m}+\alpha_{p})\delta$, and since $\frac{1}{2}\alpha_{m}\geq\alpha_{p}$, this is non-positive.
We prove part 2 now. Moving $\delta$ from $p$ to a subset of the steep projects decreases the utility of $i$ by $\alpha_{p}\delta$ and increases it by at least $0.5\alpha_{m}\delta$. Since $\frac{1}{2}\alpha_{m}>\alpha_{p}$, the sum of these is (strictly positive). If this move suppresses $j$, then the increase is more than $0.5\alpha_{m}\delta$, thus requiring $\frac{1}{2}\alpha_{m}\geq\alpha_{p}$ is enough.
To prove part 3, assume that $\frac{1}{2}\alpha_{m}<\alpha_{p}$ and we can take $\delta>0$ from some of the steep projects where $i$ receives half of the value so as to keep receiving a half of the new value. Then, moving this $\delta$ to $p$ decreases $i$’s utility from the steep projects by $0.5\alpha_{m}\delta$ and its utility from $p$ increases by $\alpha_{p}\delta$. The total change is $(-0.5\alpha_{m}+\alpha_{p})\delta$, and since $\frac{1}{2}\alpha_{m}<\alpha_{p}$, this is (strictly) positive. ∎
We prove Theorem 5 now.
Proof. We first prove case 1. Consider any NE. By Lemma 2, each player contributes to all the steep projects and if it contributes to a non-steep project, then it is the only contributor there. Take any non-steep project $p$, where someone, say player $i$, contributes a positive amount $y$. If $\alpha_{p}<0.5\alpha_{m}$, consider moving all what player $i$ contributes to $p$ to a steep project. According to part 2 of Lemma 4, this move is profitable, contradicting the assumption of an equilibrium. Therefore, $\alpha_{p}=0.5\alpha_{m}$. If the other player $j$ could move $\theta y$ to project $p$ from a steep project without losing its half at the steep project, then $j$ would strictly profit from this, because
$\frac{\alpha_{m}}{2}(\theta y)<\frac{1}{2}\frac{\alpha}{2}((\theta+1)y)\iff\theta<\frac{1}{2}(\theta+1)\iff\theta<1,$
which is always the case. Since we assume an NE, this cannot happen, implying that at any steep project, the part above the minimum required to get its half is strictly less than $\theta y$. Therefore, $i$ can move $y$ to a steep project and suppress $j$ there. Then, according to part 2 of Lemma 4, this move is profitable, contradicting the NE. Therefore, there can be no contribution to $p$. This means that only the steep projects obtain contributions, and therefore, $\operatorname{PoA}=1$. We have fully proven case 1.
Consider case 2a now. From the proof of the existence on an NE in this case, we know that $1$ investing all its budget in $m-k$ and $2$ dividing its budget equally between the steep projects constitute an NE. Thus, $\operatorname{PoS}\geq\frac{\alpha_{m}B_{2}+\alpha_{m-k}B_{1}}{\alpha_{m}(B_{1}+B_{2})}$. Since $\frac{\alpha_{m-k}}{\alpha_{m}}\leq\frac{1}{1+\theta}$, not all the projects have equal value functions. Therefore, in an NE no player is suppressed, since if $2$ dominated $1$, then $2$ would have to invest more than $B_{1}/\theta$ in each project, and $2$ would like to deviate to contribute to the steep projects more. Since no player is suppressed, we conclude from part 1 of Lemma 3 that player $1$ never contributes to a steep project in an NE, and thus $\operatorname{PoS}=\frac{\alpha_{m}B_{2}+\alpha_{m-k}B_{1}}{\alpha_{m}(B_{1}+B_{2})}$.
Next, let us approach the price of anarchy. According to Lemma 3, the only way to reduce the efficiency relatively to the price of stability is for player $2$ to invest a $2$-steep project. If this happens, then we obtain that case 2b must hold, exactly as it is done in the proof of the other direction of Theorem 3. Therefore, if this case does not hold, then $\operatorname{PoS}=\operatorname{PoS}$. If it does, then we have the NE when $2$ invests $\theta B_{1}$ in project $m-k$ (and $1$ invests all its budget there, and $2$ equally divides the rest of its budget between the steep projects), which yields the price of anarchy of $\frac{\alpha_{m}(B_{2}-\theta B_{1})+\alpha_{m-k}(B_{1}(1+\theta))}{\alpha_{m}(B_{1}+B_{2})}$.
We prove case 2b now. We show in the proof of case 2b of Theorem 3 that player $1$ investing all its budget in $m-k$ and $2$ investing $\theta B_{1}$ in $m-k$ and uniformly dividing the rest between the steep projects is an NE. Thus, $\operatorname{PoS}\geq\frac{\alpha_{m}(B_{2}-\theta B_{1})+\alpha_{m-k}(B_{1}(1+\theta))}{\alpha_{m}(B_{1}+B_{2})}$. Since $\frac{\alpha_{m-k}}{\alpha_{m}}\leq\frac{2(1-\theta)}{2-\theta}$, we conclude analogously to what we did in the proof of the previous case that no player is suppressed. Thus, Lemma 3 implies that the only way to achieve a more efficient NE is for $2$ to contribute only to the steep projects, while player $1$ contributes only to the two-steep projects. If this is an NE, then we obtain that case 2a must hold, exactly as it is done in the proof of the other direction of Theorem 3. Therefore, if this does not hold, we have $\operatorname{PoS}=\frac{\alpha_{m}(B_{2}-\theta B_{1})+\alpha_{m-k}(B_{1}(1+\theta))}{\alpha_{m}(B_{1}+B_{2})}$. If case 2a does hold, then we know that the profile where $1$ invests all its budget in $m-k$ and $2$ divides its budget between the steep projects is an NE, and thus $\operatorname{PoS}=\frac{\alpha_{m}B_{2}+\alpha_{m-k}B_{1}}{\alpha_{m}(B_{1}+B_{2})}$.
We turn to the price of anarchy now. According to Lemma 3, the NE with player $1$ investing all its budget in $m-k$ and $2$ investing $\theta B_{1}$ in $m-k$ and uniformly dividing the rest of $b_{2}$ between the steep projects is the worst possible NE, and thus $\operatorname{PoA}=\frac{\alpha_{m}(B_{2}-\theta B_{1})+\alpha_{m-k}(B_{1}(1+\theta))}{\alpha_{m}(B_{1}+B_{2})}$.
Finally, in the case 2c we know that $2$ dividing its budget equally and $1$ contributing all its budget is an NE, and therefore $\operatorname{PoS}=1$. To find the price of anarchy, recall that if $2$ does as before while $1$ invests nothing at all, it still is an NE, and thus $\operatorname{PoA}\leq\frac{\alpha_{m}B_{2}}{\alpha_{m}(B_{1}+B_{2})}$. Since $2$ always gets at least $\alpha_{m}B_{2}$ in any NE, the price of anarchy cannot decrease below it, and thus $\operatorname{PoA}=\frac{\alpha_{m}B_{2}}{\alpha_{m}(B_{1}+B_{2})}$. ∎
Let us now prove Corollary 2.
Proof. The maxima are attained in case 1 of Theorem 5.
To find the infima, find the infimum in every case, substituting the extreme values in the expressions for $\operatorname{PoS}$ and $\operatorname{PoA}$. We begin with the $\operatorname{PoS}$. In case 2a, the infimum of the $\operatorname{PoS}$ is $\frac{k}{k+\theta}$, attained for $\alpha_{m-k}=0$ and $B_{1}=\frac{\theta B_{2}}{k}$. In case 2b, the infimum of the $\operatorname{PoS}$ is the minimum of these two expressions
  1. 1.
    $\frac{\alpha_{m}B_{2}+\alpha_{m-k}B_{1}}{\alpha_{m}(B_{1}+B_{2})}$ when $\alpha_{m-k}=\frac{2\theta}{1+\theta}\alpha_{m}$ and $B_{1}=\frac{\theta B_{2}}{k+\theta^{2}}$, which is $\frac{k+\theta^{2}+\frac{2\theta^{2}}{1+\theta}}{k+\theta+\theta^{2}}$.
  2. 2.
    $\frac{\alpha_{m}(B_{2}-\theta B_{1})+\alpha_{m-k}(B_{1}(1+\theta))}{\alpha_{m}(B_{1}+B_{2})}$ when $\alpha_{m-k}=\frac{2\theta}{1+\theta}\alpha_{m}$ and $B_{1}=\frac{\theta B_{2}}{k+\theta^{2}}$, which is $\frac{k+2\theta^{2}}{k+\theta+\theta^{2}}$.
The minimum of these expressions is $\frac{k+2\theta^{2}}{k+\theta+\theta^{2}}$. Finally, the infimum of the price of stability in case 2c is $1$. The absolute infimum is the minimum of these three expressions, which is $\frac{k}{k+\theta}$.
We consider now the infimum of the price of anarchy. In case 2a, the infimum of the $\operatorname{PoA}$ is attained as follows:
  1. 1.
    If also case 2b holds, then it is the value of $\frac{\alpha_{m}(B_{2}-\theta B_{1})+\alpha_{m-k}(B_{1}(1+\theta))}{\alpha_{m}(B_{1}+B_{2})}$ when $\alpha_{m-k}=\frac{2\theta}{1+\theta}\alpha_{m}$ and $B_{1}=\frac{\theta B_{2}}{k+\theta^{2}}$, which is $\frac{k+2\theta^{2}}{k+\theta+\theta^{2}}$.
  2. 2.
    Otherwise, $\operatorname{PoA}=\operatorname{PoS}$, and so the infimum is $\frac{k}{k+\theta}$.
The minimum of these two expressions is $\frac{k}{k+\theta}$. In case 2b, the infimum of the $\operatorname{PoA}$ is $\frac{\alpha_{m}(B_{2}-\theta B_{1})+\alpha_{m-k}(B_{1}(1+\theta))}{\alpha_{m}(B_{1}+B_{2})}$ when $\alpha_{m-k}=\frac{2\theta}{1+\theta}\alpha_{m}$ and $B_{1}=\frac{\theta B_{2}}{k+\theta^{2}}$, which is $\frac{k+2\theta^{2}}{k+\theta+\theta^{2}}$. In case 2c, it is $\frac{B_{2}}{B_{1}+B_{2}}$ when $B_{1}=\frac{\theta}{m}B_{2}$, which is $\frac{m}{m+\theta}$. Therefore, the infimum of the price of anarchy is $\frac{k}{k+\theta}$. ∎
We now present the proof of Theorem 6.
Proof. We first prove case 1. According to proof of case 1 in Theorem 4, equally dividing all the budgets between the steep projects is an NE. Therefore, $\operatorname{PoS}=1$.
In order to bound the price of anarchy, we notice that since player $n$ can always obtain at least
$\frac{\alpha_{m}B_{n}(1+(n-1)\theta)}{n}$
by investing all her budget in a steep project, this is a lower bound on what she obtains in any NE. Since $B_{n-1}\geq\theta B_{n}$, we analogically conclude that player $n-1$ receives at least
$\frac{\alpha_{m}B_{n-1}(1+(n-1)\theta)}{n}$
in any Nash equilibrium. Therefore, the social welfare in an NE is at least
$\frac{\alpha_{m}(B_{n-1}+B_{n})(1+(n-1)\theta)}{n},$
which, in turn, implies the lower bound.
We consider case 2 now. The lower bound on the price of anarchy stems from the fact that player $n$ always receives at least $\alpha_{m}B_{n}$ in any NE, by investing $B_{n}$ in a steep project.
Assume now that $B_{n-1}<\frac{\theta}{\left|\Omega\right|}B_{n}$ and all the project functions are equal. We know that $n$ dividing its budget equally and all the other players contributing all their budgets is an NE, and therefore $\operatorname{PoS}=1$. To find the price of anarchy, recall that if $n$ does as before while all the other players invest nothing at all, it still is an NE, and thus $\operatorname{PoA}\leq\frac{\alpha_{m}B_{n}}{\alpha_{m}(\sum_{i\in\left\{1,2,\ldots,n\right\}}{B_{i}})}$. Since $n$ always gets at least $\alpha_{m}B_{n}$ in any NE, the price of anarchy cannot decrease below it, and thus $\operatorname{PoA}=\frac{\alpha_{m}B_{n}}{\alpha_{m}(\sum_{i\in\left\{1,2,\ldots,n\right\}}{B_{i}})}$. ∎
We now prove Corollary 3.
Proof. Theorem 7 implies the lower bound of
$\frac{1+(n-1)\theta}{n}\frac{\sum_{i=l}^{n-1}{B_{i}}}{\sum_{i=1}^{n}{B_{i}}}+\frac{1+(n-l)\theta}{n-l+1}\frac{B_{n}}{\sum_{i=1}^{n}{B_{i}}}$
on the price of anarchy. Since $\theta\leq 1$, this is at least
$\frac{\theta+(n-1)\theta}{n}\frac{\sum_{i=l}^{n-1}{B_{i}}}{\sum_{i=1}^{n}{B_{i}}}+\frac{\theta+(n-l)\theta}{n-l+1}\frac{B_{n}}{\sum_{i=1}^{n}{B_{i}}}=\theta\frac{\sum_{i=l}^{n-1}{B_{i}}}{\sum_{i=1}^{n}{B_{i}}}+\theta\frac{B_{n}}{\sum_{i=1}^{n}{B_{i}}}=\theta\frac{\sum_{i=l}^{n}{B_{i}}}{\sum_{i=1}^{n}{B_{i}}},$
providing the lower bound. ∎
We shall now prove Theorem 9. To remain self-contained, before proving the theorem, we bring here the necessary definitions used by [32]. Given player $i$ with the strategy set $A_{i}\subseteq\mathbb{R}^{m}$, define $A\stackrel{{\scriptstyle\rm\Delta}}{{=}}A_{1}\times\ldots\times A_{n}$.
Definition 5.
For each pair of players $i,j\in{1,\ldots,n}$, let $D(i)$ be a positive natural, and for a $d\in\left\{1,\ldots,D(i)\right\}$, let $f^{d}_{i,j}\colon\mathbb{R}\to\mathbb{R}$ be continuous, such that $(f^{d}_{i,j})^{-1}=f^{d}_{j,i}$. For every player $i$, we define
$A^{*}(i)\stackrel{{\scriptstyle\rm\Delta}}{{=}}\{(a_{1},\ldots,a_{n})\in A\mid\exists j\neq i,\exists k\in\left\{1,\ldots,m\right\},\exists d\in\left\{1,\ldots,D(i)\right\},$
[…]
We now define weakly lower semi-continuity, which intuitively means that there is a set of directions, such that approaching a point from any of these directions gives values at least equal to the function at the point.
Definition 6.
Let $B^{m}\stackrel{{\scriptstyle\rm\Delta}}{{=}}\left\{z\in\mathbb{R}^{m}\mid\sum_{i=1}^{m}{{z_{l}}^{2}}=1\right\}$, i.e. the surface of the unit sphere centered at zero. Let $e\in B^{m}$ and $\theta>0$. Function $g_{i}(a_{i},a_{-i})$ is weakly lower semi-continuous in the coordinates of $a_{i}$ if for all $\hat{a_{i}}$ there exists an absolutely continuous measure $\nu$ on $B^{m}$, such that for all $a_{-i}$, we have
$\int_{B^{m}}\left\{\liminf_{\theta\to 0}g_{i}(\hat{a_{i}}+\theta e,a_{-i})d\nu(e)\right\}\geq g_{i}(\hat{a_{i}},a_{-i}).$
Finally, we are ready to prove Theorem 9.
Proof. We show now that all the conditions of Theorem $5^{*}$ from [32] hold. First, the strategy set of player $i$ is simplex, and as such, it is non-empty, convex and compact. The utility function of player $i$ is discontinuous only at a threshold of one of the projects. These points belong to the set $A^{*}(i)$, defined in Formula(s) (3), if we take
$D(i)\stackrel{{\scriptstyle\rm\Delta}}{{=}}2;$
$f^{1}_{i,j}(y)\stackrel{{\scriptstyle\rm\Delta}}{{=}}y\begin{cases}\theta&\text{if }i<j,\\ 1/\theta&\text{if }i>j;\end{cases}$
$f^{2}_{i,j}(y)\stackrel{{\scriptstyle\rm\Delta}}{{=}}y\begin{cases}\theta&\text{if }j<i,\\ 1/\theta&\text{if }j>i.\end{cases}$
The sum of all the utilities is a continuous function. In addition, the utility of player $i$ is bounded by the largest project’s value when all the players contribute their budgets there. It is also weakly lower semi-continuous in $i$’s contribution, since if we take the measure $\nu$ to be $\nu(S)\stackrel{{\scriptstyle\rm\Delta}}{{=}}\lambda(S\cap B^{m}+)$, where $\lambda$ is the Lebesgue measure on $B^{m}$ and $B^{m}+\stackrel{{\scriptstyle\rm\Delta}}{{=}}\left\{z\in B^{m}\mid z_{l}\geq 0,\forall l=1,\ldots,m\right\}$, we obtain an absolutely continuous measure $\nu$, such that the integral sums up only the convergences to a point from the positive directions, and such convergences will never become less than the function at the point.
Finally, Theorem $5^{*}$ implies that a mixed NE exists. ∎
Appendix B Simulations
The theory provides only sufficiency results for shared effort games with more than $2$ players. To better explore the existence and efficiency of NE in these games, we simulate a variation of fictitious play [45]; intuitively, we try fictitious play because of its well-known properties of converging to an NE in some finite cases (see Section 6). The main value of this section is in the methodology; the concrete simulations are best done based on the studied case at hand.
We first adapt the fictitious play [12] to our infinite game. Danskin [14] defines the best response to maximise the average utility against all the previous strategies of the other players, while we, as well as the original fictitious play, best respond to the averaged (cumulative) strategy of the others.
Aspiring to implement the adapted fictitious play, we then suggest an algorithm for finding a best response, if it exists, for the case of two projects. For more projects, we prove that even best responding is already NP-hard.
B.1 Infinite-Strategy Fictitious Play for Shared Effort Games
In fictitious play, a player best responds to the averaged strategies of the other players. Since the game has convex strategy spaces (simplexes), we do not need mixing to average the strategies. Each player’s strategy is averaged separately, and each player best responds to the product distribution of the other players. Denote the set of all the best responses of player $i$ to profile $x^{-i}$ of the others by $\operatorname{BR}(x^{-i})$.
Definition 7.
Given a shared effort game with players $N$, budget-defined strategies $S^{i}=\left\{x^{i}=(x^{i}_{\omega})_{\omega\in\Omega}\in\mathbb{R}_{+}^{\left|\Omega\right|}\mid\sum_{\omega\in\Omega}{x^{i}_{\omega}}\leq B_{i}\right\}$ and utilities $u^{i}(x)\stackrel{{\scriptstyle\rm\Delta}}{{=}}\sum_{\omega\in\Omega}{\phi^{i}_{\omega}(x_{\omega})}$, define an Infinite-Strategy Fictitious Play (ISFP) as the following set of sequences. Consider a (pure) strategy in this game at time $1$, i.e. $(x^{i}(1))^{i\in N}=((x^{i}(1)_{\omega})_{\omega\in\Omega})^{i\in N}$, define $X^{i}(1)\stackrel{{\scriptstyle\rm\Delta}}{{=}}\left\{x^{i}(1)\right\}$, and define recursively, for each $i\in N$ and $t\geq 0$, the set of the possible strategic histories at time $t+1$ to be all the possible combinations of the history with the best responses to the previous strategic histories:
$X^{i}(t+1)\stackrel{{\scriptstyle\rm\Delta}}{{=}}\left\{\frac{tx^{i}(t)+\operatorname{br}(x^{-i}(t))}{t+1}\mid x^{i}(t)\in X^{i}(t),\operatorname{br}(x^{-i}(t))\in\operatorname{BR}(x^{-i}(t))\right\}.$ (4)
We say that an ISFP converges to $x^{*}\in{\mathbb{R}_{+}}^{n}$ if at least one of its sequences converges to $x^{*}$ in every coordinate. Please note that if at a given time the sequence plays an NE, it can stay there forever.
Since $\operatorname{BR}(x^{-i}(t))$ is a set, there may be multiple ISFP sequences. For an ISFP to be defined, we need that $\operatorname{BR}(x^{-i}(t))\neq\emptyset$, meaning that the utility functions attain a maximum. However, the functions are, generally speaking, not upper semi-continuous, and may sometimes not attain a maximum.
In ISFP, all the plays have equal weights in the averaging. In the other extreme, a player just best-responds to the previous strategy profile of other players, thereby attributing the last play with the weight of $1$ and all the other plays with 00. In general, we define, for an $\alpha\in[0,\infty]$, an $\alpha-\text{ISFP}$ play as in Definition 7, but with the following formula instead [^16] For $\alpha=\infty$, we just obtain a constant sequence.
$X^{i}(t+1)\stackrel{{\scriptstyle\rm\Delta}}{{=}}\left\{\frac{\alpha tx^{i}(t)+\operatorname{br}(x^{-i}(t))}{\alpha t+1}\mid x^{i}(t)\in X^{i}(t),\operatorname{br}(x^{-i}(t))\in\operatorname{BR}(x^{-i}(t))\right\}.$ (5)
Here, the last play’s weight is $\frac{1}{\alpha t+1}$.
We do not know whether and when any convergence property can be proven for the generalised fictitious play in shared effort games. In simulations, our generalised fictitious play often converges to an NE, bolstering the importance of equilibrium efficiency, as stated by Roughgarden and Tardos [39, Chapter 17].
Next, we solve the algorithmic problem of finding whether a best response exists, and if it does, what it is.
B.2 Best Response
First, we show that finding a best response is NP-hard, even when the game has only two players, and then, we do find it in polynomial time for a $2$-project game.
Theorem 10.
Best-responding (if a best response exists and otherwise returning that none exists) in shared effort games with $\theta$-equal sharing is NP-hard already for two players, even if all the project functions have the same coefficients.
Proof. We reduce the following NP-hard problem [31, Section 1.3] to best-responding by, w.l.o.g., player $1$.
Definition 8.
The subset sum problem receives items $\left\{1,\ldots,n\right\}$ of sizes $s_{1},\ldots,s_{n}$ and a cap $C$, and returns a subset $S$ of $\left\{s_{1},\ldots,s_{n}\right\}$ such that $\sum_{i\in S}{s_{i}}$ is maximum possible that is at most $C$.
The reduction proceeds as follows. Let $B_{1}$ be $C$. For each item $i$, create project $i$, where player $2$ contributes $s_{i}/\theta$, where $\theta$ is set small enough so that player $1$ will not be able to suppress $2$, i.e., $C\leq s_{i}/(\theta^{2})$, for each $i$. First, the inability to suppress makes player $1$’s utility upper semi-continuous, so a best response exists. All the equal project coefficients $\alpha$ are such that $\alpha C<s_{i},\forall i=1,\ldots,n$, so that it is always better to achieve the threshold at yet another project than to contribute more to a project where the threshold is already achieved. We finally require that $\alpha(1+\theta)=2\theta$, which can be achieved by making $\alpha$ and/or $\theta$ smaller.
Fix any best response in the obtained game. The projects where $1$ contributes above the threshold constitute an optimal solution to the subset sum problem, because player $1$ needs to provide at least $\theta(s_{i}/\theta)=s_{i}$ in order to be above the threshold at project $i$, and such a contribution provides it with the utility of $\frac{\alpha(s_{i}/\theta+s_{i})}{2}$, which is, because of the requirement $\alpha(1+\theta)=2\theta$, equal to $s_{i}$. The equivalence is completed by the fact that contributing more to such a project is less important than achieving the threshold at another project, because $\alpha C<s_{i}$. ∎
In order to execute the infinite-strategy fictitious play, from now on we assume that the game has only two projects, $\Omega=\left\{\psi,\omega\right\}$. We would like to find a best response for a player $i\in N$, all the other players’ strategies $x^{-i}\in S^{-i}$ being fixed. From the weak monotonicity of the sharing functions $\phi^{i}_{\omega}$, we may assume w.l.o.g. that a best responding player contributes all her budget. Then, a strategy is uniquely determined by the contribution to project $\psi$ and we shall write $x^{i}$ for $x^{i}_{\psi}$, meaning that $x^{i}_{\omega}=B_{i}-x^{i}$.
We now state the conditions for the following theorem. Consider $M_{\text{eq}}^{\theta}$ sharing and convex project value functions. Let $D^{i}_{0}<D^{i}_{1}<\ldots<D^{i}_{m}$ and $W^{i}_{0}<W^{i}_{1}<\ldots<W^{i}_{l}$ be the jumps of $\phi^{i}_{\psi}$ and $\phi^{i}_{\omega}$ respectively. (The first points in each list are the minimum contributions to projects $\psi,\omega$, respectively, required for $i$ to obtain a share. The other points are the points at which another player becomes suppressed at the respective project.) The possible discontinuity points of the total utility of $i$ are thus $D^{i}_{0}<D^{i}_{1}<\ldots<D^{i}_{m}$ and $B_{i}-W^{i}_{l}<\ldots<B_{i}-W^{i}_{1}<B_{i}-W^{i}_{0}$. Denote the distinct points of these lists merged in the increasing order by $L$. Let $L_{B_{i}}$ denote the points of the list $L$ that are on $[0,B_{i}]$, together with 00 and $B_{i}$, and let $M_{B_{i}}$ be $L_{B_{i}}$ with an arbitrary point added between each two consecutive points.
Theorem 11.
The maximum of the one-sided limits at the points of $L_{B_{i}}$ and of the values at the points of $M_{B_{i}}$ yields the utility supremum [^17] The supremum is the exact upper bound; it always exists. of the responses of player $i$. This supremum is a maximum (and in particular, a best response exists) if and only if it is achieved at a point of $M_{B_{i}}$.
Proof. The utility of $i$ is $u^{i}(x^{i})=\phi^{i}_{\psi}(x^{i})+\phi^{i}_{\omega}(B_{i}-x^{i})$. Consider the open intervals between the consecutive points of $L_{B_{i}}$. On each of these intervals, the function $\phi^{i}_{\psi}(x^{i})$ is convex, being proportional to the convex project value function, and $\phi^{i}_{\omega}(B_{i}-x^{i})$ is convex because the function $B_{i}-x$ is convex and concave and $\phi^{i}_{\omega}$ is convex and weakly monotone. Therefore, the utility is also convex, as the sum of convex functions.
Therefore, the supremum of the utility on the closure of such a convexity interval is attained as the one-sided limit of at least one of its edge points. This supremum can be a maximum if and only if it is not larger than the maximum of the utility at an interval edge point or at an internal point of an interval (in the last case, the convexity implies that the utility is constant on this interval). ∎
When finding the one-sided limit at a point of $L_{B_{i}}$ takes constant time, the resultant algorithm runs in $O(n\log n)$ time and in linear space. We employ this algorithm in the simulations.
B.3 The Simulation Method
[…]
We choose an initial belief state about all the players and run the ISFP from this state on, updating this common belief state at each step by finding a best response of each player to the current belief state and averaging it with the history. To increase the chances of finding an existing NE, for each game, we generate $45$ fictitious plays by randomly and independently generating the initial belief state on each player’s actions, uniformly over the possible histories (this number was experimentally found to be a good balance between run time and precision). While simulating, when a player has multiple best responses, we choose a closest one to the current belief state of the fictitious play, in the sense of minimizing the maximum distance from the last action’s components. For each found NE, we calculate its efficiency by dividing its total utility by the optimum possible total utility and plot it using shades of gray. When no NE is found, we plot it black.
B.4 Results and Conclusions
Figure 2: The existence and efficiency of NE for $2$ players as a function of the ratio of the project functions coefficients and the ratio of the two largest budgets. The first row plots the results of the simulations, and the second row shows the theoretical predictions. Black means that an equilibrium has not been found.
We present the results representing the simulation trends; We have obtained similar results in more extensive simulations as well.
First, to validate our simulations, we compare the results of the simulations to the theoretical predictions in Figure 2, where we go over the ratios of the project coefficients and budgets, because the invariance to multiplication from Proposition 2 implies that only the ratios matter for NE. This is not exactly the case here, since the budgets are equally distanced, and multiplying the largest budgets by the same factor may multiply the other budgets differently. However, ratios constitute a good start. In this case, the simulations are in complete agreement with the theory.
The difference between the areas in Figure 2 that correspond to $B_{1}\geq\theta B_{2}$ and to $B_{1}<\theta B_{2}$ fits Theorem 3. For two players, except for the NE when the two projects are equal, an NE exists for a budget ratio if and only if the value functions ratio is below a certain value, fitting Corollary 1.
Figure 3: The existence and efficiency of NE as function of project functions for $2,3,4,5$ and $6$ players. Black means that Nash Equilibrium has not been found.
Figure 4: The efficiency as a function of the ratio of the project value functions for $2,3,4,5$ and $6$ players. Efficiency of 00 means that Nash Equilibrium has not been found.
When the project function coefficients are the independent variables, Figure 3 presents the NE. Each line of the simulation plots corresponds to a setting for a given number of players, and within a line, the plots are generated for an increasing sequence of $\theta$. Mostly, an NE exists except for a cone where the project functions are quite close to each other. Interestingly, sometimes an NE exists also when the project functions are nearly the same (at ratio $1$), or at another constant ratio with each other. In all cases, the ratio of project functions determines existence of an NE, as Proposition 2 implies. Usually, the more players there are, the fewer settings with an NE we find.
Figure 4 demonstrates efficiency as a function of the ratio of the coefficients of the project functions. This efficiency is uniquely determined by the ratio of the project functions, complying with Proposition 2. The dependency is piecewise linear and non-decreasing in each linear interval. For two players, it is linear, in the spirit of Theorem 5. (Though not directly predicted by it, since the theorem considers extremely efficient or inefficient equilibria, and has several cases, which imply piecewise linearity.) The larger the $\theta$ is, the steeper the piecewise linear dependency becomes.
Figure 5: The existence and efficiency of NE as function of the largest and the second largest budgets. Black means that Nash Equilibrium has not been found, and gray hatching indicates the non-defined area, since the second highest budget may not be larger than the highest one.
Figure 6: The efficiency as a function of the ratio of the two largest budgets. Efficiency of 00 means that Nash Equilibrium has not been found.
When the largest and the second largest budgets are the independent variables, Figure 5 presents the equilibria. (The budgets of the other players are spread on equal intervals). Each line of the plot corresponds to a setting of project function coefficients for a given number of players, and within a line, the plots are generated for an increasing sequence of $\theta$. The more players we have, the fewer NE we typically find. Figure 6 plots efficiency as a function of the ratio of the two largest budgets.
To summarise, the existence of NE is related to the ratio of project function coefficients and the budget ratio being in some limits, limits that in particular depend on the threshold. This is in the spirit of Theorem 2, Proposition 2 and Theorems 3 and 4. Based on Theorem 2, Proposition 2, Theorem 3, Corollary 1, and the simulation results, we hypothesise that also for multiple projects and players, an NE exists if and only if at least one of several sets of conditions on the ratios of the budgets holds, and for every such set, several conditions of being smaller or equal or exactly equal to a function of the threshold (and not the budgets) on the ratios of project function coefficients hold together.
The efficiency of the NE that we find depends on the ratio of the project function coefficients and on the ratio of the budgets, rather than on the project functions and budgets themselves, in the spirit of Proposition 1, Proposition 2. Based on Proposition 1, Proposition 2, Theorem 5 and the simulation results, we hypothesise that also for multiple projects and players, the price of anarchy and stability of a shared effort game depends piecewise linearly on the project function coefficients ratios, and within each linear domain, this dependency is non-decreasing with $\theta$.